### v1

# Adaptive Voxel-Weighted Loss Using L1 Norms in Deep Neural Networks for Detection and Segmentation of Prostate Cancer Lesions in PET/CT Images

Abstract
This study proposes a new loss function for deep neural networks, L1-weighted Dice Focal Loss (L1DFL), that leverages L1 norms for adaptive weighting of voxels based on their classification difficulty, towards automated detection and segmentation of metastatic prostate cancer lesions in PET/CT scans. We obtained 380 prostate-specific membrane antigen (PSMA) [18F]DCFPyL PET/CT scans of patients diagnosed with biochemical recurrence metastatic prostate cancer. We trained two deep neural networks with 3D convolutional neural network backbone — Attention U-Net and SegResNet — and concatenated the PET and CT volumes channel-wise as input. The performance of our custom loss function was evaluated against the Dice and Dice Focal Loss functions. For clinical significance, we considered a detected region of interest (ROI) as a true positive if at least the voxel with the maximum standardized uptake value falls within the ROI. We assessed the models’ performance based on the number of lesions in an image, tumor volume, activity and extent of lesion spread. The L1DFL outperformed the comparative loss functions by at least $13\%$ on the test set. In addition, the F1 scores of the Dice Loss and the Dice Focal Loss were lower than that of L1DFL by at least $6\%$ and $34\%$, respectively. The Dice Focal Loss yielded more false positives, whereas the Dice Loss was more sensitive to smaller volumes and struggled to segment larger lesions accurately. They also exhibited network-specific variations and yielded declines in segmentation accuracy with increased tumor spread. Our results demonstrate the potential of L1DFL to yield robust segmentation of metastatic prostate cancer lesions in PSMA PET/CT images. The results further highlight potential complexities arising from the variations in lesion characteristics that may influence automated prostate cancer tumor detection and segmentation. The code is publicly available at: https://github.com/ObedDzik/pca_segment.git.
keywords:
Detection , segmentation , L1 norm, adaptive weighting, loss function, PSMA PET/CT imaging
[…]
[a]organization=School of Biomedical Engineering,addressline=University of British Columbia, city=Vancouver, country=Canada
[…]
[b]organization=Department of Integrative Oncology,addressline=BC Cancer Research Institute, city=Vancouver, country=Canada
[…]
[c]organization=Vector Institute,city=Toronto, country=Canada
[…]
[d]organization=Department of Physics & Astronomy,addressline=University of British Columbia, city=Vancouver, country=Canada
[…]
[e]organization=Department of Radiology,addressline=University of British Columbia, city=Vancouver, country=Canada
[…]
[f]organization=Department of Electrical and Computer Engineering,addressline=University of British Columbia, city=Vancouver, country=Canada
1 Introduction
Semantic segmentation involves mapping every pixel of an image to a class, whether an object or a background [1]. It requires both a local and contextual understanding of the image [2], making it particularly difficult since neighboring pixels are highly related; thus, information on the spatial context and the boundary’s smoothness is essential in a segmentation model [3]. Particularly in medical image analysis, higher accuracy levels are required for tumor segmentation since false detections may significantly impact patient management [4]. However, medical images are fraught with an imbalance in the proportion of voxels belonging to the regions of interest (ROIs) compared to the background making a balance in precision and sensitivity in automated segmentation a challenging task.
Whole body positron emission tomography (PET) scans of cancer patients are an example of such medical images where voxels corresponding to the foreground class are overwhelmingly outnumbered by background voxels. These images are characterized by high variability in lesion size, shape, and intensity, as well as significant heterogeneity in tracer uptake [5]. In the case of metastatic prostate cancer (mPCa) images, lesions can range from small, low-contrast regions barely distinguishable from surrounding tissues to large, high-uptake areas with well-defined boundaries, and may occupy different anatomical sites [6]. In localizing these lesions, the prostate-specific membrane antigen (PSMA) is often targeted due to its high expression by tumor cells. However, PSMA expression is not exclusive to prostate cancer lesions; it is also found in other tissues, including salivary glands, kidneys, and the small intestine, as well as in certain non-prostatic malignancies and inflammatory processes [7]. This broad expression can lead to radiotracer uptake in non-cancerous or non-prostatic regions challenging the specificity of computer-aided detection algorithms. Thus, the challenge in artificial intelligence (AI) based lesion detection and segmentation in PET scans of metastatic prostate cancer lesions is avoiding false-positive segmentation of anatomical structures and other artifacts while maintaining accurate lesion detection.
With most of the automated approaches, improved sensitivity of mPCa lesions often comes at the cost of specificity, yet both extremes pose potential risks. One commonly used strategy for limiting false positive rate is standardized uptake value (SUV) thresholding, which involves defining all voxels with SUV greater than some value, typically considering the background uptake of the liver to be foreground and all other voxels as background [8]. This method may be included as part of the automated segmentation process or as a post-processing step and has been shown to reduce false positive counts [9]. Yet, its major drawback is the potential of having a reduced sensitivity for lesions that exhibit lower avidity for the radiotracer, or in patients with extensive liver metastases [10].
Loss function design is another strategy that has received much attention and has led to an improvement in the performance of AI-based segmentation methods. The widely used loss function is the Dice Loss. It is based on the Dice Similarity Coefficient (DSC) and penalizes pixel-wise discrepancies between the predicted maps and their corresponding ground truth for each class [11]. A model trained using the Dice Loss objective maximizes DSC inherently incorporating a balance in sensitivity and specificity [12]. Nevertheless, the Dice Loss is not well-suited for handling another type of imbalance caused by a difference in the difficulty in classifying pixels [13]. Within a class, be it the minority foreground class or majority background class, some pixels might be easier to classify, with models achieving greater than $50\%$ probability of correctly classifying such pixels [13]. However, others may be more challenging to predict, yielding lower probabilities [13] and are therefore designated hard to classify. Without a focusing strategy implemented, the Dice Loss can be easily overwhelmed by the easy-to-classify samples and may perform poorly on the harder ones [14]. Besides, it exhibits an inconsistent performance when segmenting multiple lesions of varying sizes in an image [15].
These limitations have prompted the development of alternative loss functions. Distribution-based loss functions like Focal and Cross-Entropy losses are frequently used alternatives due to their dynamic focusing ability. They are often combined with the Dice Loss and its variants to form compound loss [11] since the losses based on overlap measurements are more robust for medical segmentation tasks [16]. With an additional ability to control the model’s focus on detecting ROIs, these combo losses generalize well [11, 17]. The challenge, however, remains on the balance of sensitivity and specificity as these compound losses tend to lose the Dice Loss component, potentially leading to more false positive segmentation [18]. Thus, various regularizations and additional constraints are required to improve model accuracy for a medical lesion segmentation task [19].
This work proposes a new loss function, L1-weighted Dice Focal Loss (L1DFL), with a dynamic weighting strategy based on L1 norms of predicted probabilities and ground truth labels. Using L1 norms, the loss function considers the difficulty of each sample and the prevalence of similar samples in the dataset. In medical image segmentation tasks, datasets are often imbalanced (disproportionately higher number of background voxels than ROIs) and dominated by easy samples (voxels that are easily classified correctly); the L1 norms of these easy samples tend to approach zero. Our technique focuses on down-weighting both easy samples and those with high L1 norm counts, which are likely to represent the background. We demonstrate the efficiency of L1DFL by experiments on [18F]DCFPyL PET/CT scans of patients with biochemical recurrence metastatic prostate cancer, for which lesions have varied sizes, primarily small, with low uptake values.
The main contributions of this work are as follows: (i) we introduce a novel loss function, L1-weighted Dice Focal Loss (L1DFL), incorporating a dynamic weighting strategy; (ii) we evaluate the performance of L1DFL, Dice Loss, and Dice Focal Loss functions in handling false positives and false negative rates; (iii) we compare the segmentation performance of the loss functions based on different lesion scenarios: single-lesion, and multiple lesions. These scenarios are based on the number of lesions in an image; (iv) we evaluate the performance of these loss functions on different clinical metrics like molecular tumor volumes and the extent of lesion spread in the body; and (v) we assess performance using two network architectures, Attention U-Net [20] and SegResNet [21].
2 Materials and Methods
2.1 Dataset
We analyzed 380 prostate-specific membrane antigen (PSMA) [18F]DCFPyL PET/CT scans of patients diagnosed with biochemical recurrence metastatic prostate cancer as part of an ongoing clinical trial (NCT02899312) with informed written consent obtained from the subjects. The image data collection was approved by the University of British Columbia – BC Cancer Ethics Board. The inclusion criteria for the patients were: (1) histologically proven prostate cancer with biochemical recurrence after initial curative therapy with radical prostatectomy, with a PSA $>$ 0.4 ng/mL and an additional PSA measurement confirming an increase; and (2) histologically proven prostate cancer with biochemical recurrence after initial curative radiotherapy, with a PSA $>$ 2 ng/mL after therapy, with five or fewer lesions identified on [18F]DCFPyL PET/CT [22].
Each patient received an average dose of 350 MBq [18F]DCFPyL after fasting for four hours, with dose adjustments based on body weight. Imaging was conducted 120 minutes after injection using a GE Discovery 600 or 690 scanner (GE Healthcare, USA). A non-contrast-enhanced CT scan was acquired for localization and attenuation correction (120 kV, automatic mA adjustment from 30–200 mA, and a noise index of 20). Following CT acquisition, a whole-body PET scan was conducted at 2–4 minutes per bed position, depending on the participant’s size, with reconstruction performed using an ordered subset expectation maximization algorithm and point-spread function modeling. Each transaxial PET image had a matrix resolution of 192 × 192 pixels, with a physical pixel size of $3.64\text{mm}^{2}$. An expert nuclear medicine physician performed manual segmentation on all active lesions [22].
There were a total of 684 prostate cancer (PCa) lesions across the entire dataset with a mean active lesion volume of 6.68 $\pm$ 10.20 ml. The lesions varied in uptake with the average maximum standardized uptake value (SUVmax) and mean standardized uptake value (SUVmean) being 12.65$\pm$14.46 and 4.62 $\pm$ 3.88, respectively.
2.2 Ground Truth Annotation
The ground truth segmentations were performed by a board-certified nuclear medicine physician from the BC Cancer Research Institute. All lesions were annotated using the PET Edge tool, a semi-automated gradient-based segmentation tool, and contours refined using the 3D Brush tool. Both tools are available in the MIM workstation (MIM software, Ohio, USA). These segmentations served as the ground truth labels for training the models.
2.3 Data Preprocessing and Augmentation
[…]
2.4 Architectures
In this work, we implemented and studied two convolutional neural network architectures: Attention U-Net [20] and SegResNet [21]. These networks, built on top of a 3D-CNN backbone, comprise encoder and decoder blocks. The SegResNet uses residual blocks with skip connections for feature propagation and maintaining gradient stability during training [21]. The encoder path contained blocks with the following number of layers: 1, 2, 2, and 4, allowing for increasing levels of feature extraction at different resolutions. The initial filter size was set to 16 channels, then doubled through each downsampling step to capture progressively higher-level features. Each convolutional layer employs a ReLU activation function, and Group Normalization [23]. The decoder path, containing one layer in each block, gradually reconstructs the segmented output while leveraging skip connections to preserve spatial details from the encoder. The input and output channels were set to 2, accommodating multi-channel input images and dual-class segmented outputs.
On the other hand, Attention U-Net extends the U-Net architecture to include attention gates along the encoder-decoder path, allowing the network to selectively pay attention to essential features [20]. The encoder path consisted of five layers, with each layer increasing in channel depth from 16, 32, 64, 128, to 256 channels, and each layer downsampled by a stride of 2. These attention gates guide the network in emphasizing regions that contribute significantly to the segmentation task. Each convolutional layer employs a ReLU activation function, and Batch Normalization as in the original implementation [20]. The output layer matched the input with two channels for multi-modal inputs and dual-class output, representing background and foreground classes.
2.5 Model Training
We divided the dataset into training, validation, and testing sets containing 258, 65, and 57 samples, respectively. The training objective was to minimize the loss functions in the training set. We used AdamW optimizer with weight decay of $10^{-5}$ to optimize the loss functions. We adopted a cosine annealing scheduler to decay the learning rate from $2\times 10^{-4}$ to zero for 1000 epochs. The loss was first computed for each batch within an epoch and the overall loss for an epoch was calculated by taking a mean over all the batch losses. The model with the highest mean DSC in the validation phase was chosen for further evaluation of the test set. The code for the implementations is made available at: https://github.com/ObedDzik/pca_segment.git
2.6 Loss Functions
Dice Loss (DL):
The Dice Loss maximizes the overlap between the predicted segmentation and the ground truth. In this work, for a specific cubic patch of an image, the Dice Loss for a batch was computed as,
$\mathcal{L}_{\text{Dice}}=1-\frac{1}{2}\sum_{c\in\{0,1\}}\frac{2\sum_{i}p_{i}(c)g_{i}(c)}{\sum_{i}p_{i}(c)+\sum_{i}g_{i}(c)+\epsilon}$ (1)
where $p_{i}$ and $g_{i}$ are, respectively, the predicted probability and ground truth for a given class $c$ of the voxel $i$, and $\epsilon$ is a small constant to prevent division by zero.
Focal Loss:
The Focal Loss [24] helps focus on hard-to-classify examples by down-weighting the easy ones. In this work, the Focal Loss for an image patch was computed over a batch as,
$\mathcal{L}_{\text{Focal}}=-\frac{1}{2}\sum_{c\in\{0,1\}}\sum_{i}\alpha_{c}(1-p_{i}(c))^{\gamma}\log(p_{i}(c))$ (2)
where $p_{i}$ is the predicted probability for a given class $c$ of a voxel $i$, $\alpha$ is a factor defining the balance between background and foreground classes in a binary segmentation task, and $\gamma$ is the focusing parameter that adjusts the rate at which easy examples are down-weighted.
Dice Focal Loss (DFL):
The Dice Focal Loss [25] combines the Dice Loss and the Focal Loss forming a compound loss function:
$\mathcal{L}_{\text{DFL}}=\mathcal{L}_{\text{Dice}}+\mathcal{L}_{\text{Focal}}$ (3)
For the Dice Focal Loss formulation, we set $\gamma=2$ and $\alpha=1$ in all our experiments similar to the implementation reported in [26].
L1-weighted Dice Focal Loss (L1DFL):
The L1-weighted Dice Focal Loss (L1DFL) applies a weighting strategy to the Dice Loss (DL) based on the L1 norm between the predicted probabilities and the ground truth labels and combines it with the Focal Loss (FL). The Dice Loss generally has two variants, one with squared terms in the denominator [27] and one without [12]. In L1DFL, we employ the Dice Loss with the squared denominator which was preferred over the non-squared version following its performance in [28]. We describe the mathematical formulation of the weighting strategy next.
First, we compute the L1 norm, $\Delta$, between the predicted probabilities $p$ and the ground truth labels $g$,
$\Delta_{i}=\|g_{i}-p_{i}\|_{1},\text{ for $i=1,2,...N$ }$ (4)
where N is the total number of samples. Next, we partition the range of L1 norm values $\Delta$ into bins of a consistent nominal bin width of $\Gamma$. Each bin $B$ is defined by its center $B_{k}$:
$B_{k}=\Gamma\cdot k,\text{for $k=0,1,2,...n-1$}$ (5)
where $n=\lceil{\frac{1}{\Gamma}}+1\rceil$ is the total number of bins. For instance, with a bin width of $\Gamma=0.1$ and a total of $n$ = 11 bins, $B_{k}$ takes values at regular intervals in the range [0,1], such that for $k=0,1,2,\dots,10$, bin centers ($B_{k}$) = $[0,0.1,0.2,\dots,1.0]$.
We then calculate the count of $\Delta$ values that fall within each bin:
$\mathcal{C}(B_{k})=\sum_{i=1}^{N}\delta_{\kappa}(B_{k},\Delta_{i})$ (6)
where $\delta_{\kappa}(B_{k},\Delta_{i})$ is an indicator function defined as:
$\delta_{\kappa}(B_{k},\Delta_{i})=\begin{cases}1,&\text{if }|B_{k}-\Delta_{i}|\leq\frac{\Gamma}{2}\\ 0,&\text{otherwise}\end{cases}$ (7)
To account for potential truncation near the boundaries of [0,1], we calculate the effective bin width for each bin as:
$\lambda_{\kappa}(B_{k})=\min\Big{\{}B_{k}+\frac{\Gamma}{2},1\Big{\}}-\max\Big{\{}B_{k}-\frac{\Gamma}{2},0\Big{\}}$ (8)
Using $\lambda_{\kappa}(B_{k})$, the norm density, $\mathcal{D}(B)$, for each bin is defined as:
$\mathcal{D}(B_{k})=\frac{C(B_{k})}{\lambda_{\kappa}(B_{k})}$ (9)
The density $\mathcal{D}(B_{k})$ represents the concentration of samples around the bin center $B_{k}$. For $B_{k}$ values close to 0, the L1 norm values ($\Delta_{i}$) are also close to 00, indicating easy samples where the predictions align closely with the ground truth. In contrast, bins with centers near 1 represent L1 norms closer to 1, indicating hard-to-classify samples where predictions deviate significantly from the ground truth. Thus, based on the norm density, we can evaluate how common or rare a particular prediction difficulty is, allowing the weighting to be adjusted accordingly. This adjustment ensures that less frequent and difficult samples, which often correspond to foreground samples in an imbalanced scenario receive higher weights than the easy-to-classify samples. The weight for each bin is then calculated as:
$w(B_{k})=\frac{N}{\mathcal{D}(B_{k})}$ (10)
These weights are applied to the numerator and the denominator of the Dice Loss to account for the imbalances caused by the variations in $\Delta$
$\mathcal{L}_{\text{wDice}}=1-\frac{2\sum_{i=1}^{N}w_{i}y_{i}p_{i}+\epsilon}{\sum_{i=1}^{N}w_{i}(y_{i}^{2}+p_{i}^{2})+\epsilon}$ (11)
where $w_{i}=w(B_{k})$ if $\Delta_{i}$ belongs to bin $B_{k}$. If the L1 norms of the examples are uniformly distributed, the density $\mathcal{D}(B_{k})$ will be the same for all bins, resulting in equal weights $w(B_{k})$ for all bins. In this case, the weighted Dice Loss, ($L_{\text{wDice}}$), will reduce to the standard Dice Loss.
Finally, we combine $\mathcal{L}_{\text{wDice}}$ with the Focal Loss. The full expression of L1DFL thus is:
[…]
We illustrate the dynamic weighting strategy of L1DFL below. For a given 4x4 matrix, Figure 1 highlights how L1DFL downweights contributions from the background class and focuses on regions of classification difficulty. Specifically, the figure shows that higher weights are applied for voxels with L1 norm ($\Delta$) close to 1, but for voxels with L1 norms close to 0, representing that those voxels are easily classified, lower weights are applied. Thus, regions corresponding to false positives and negatives receive much attention from the model. In our experiments, we empirically selected a constant bin width, $\Gamma$, of 0.1 and $\gamma$ of 2 for the focal loss component of the loss function.
Figure 1: An illustration of the dynamic weighting strategy in L1 weighted Dice Focal Loss (L1DFL). (a) shows the ground truth (G), where 1 indicates lesion areas and 0 indicates background. (b) displays the predicted probability map, where green regions denote false positive predictions and red regions represent ROIs. (c) shows the L1 norm values between the predicted probabilities and ground truth, highlighting the error magnitude for each voxel. Higher values imply higher classification difficulty. (d) presents the density, $\mathcal{D}$, which reflects the frequency distribution of each L1 norm value, capturing the difficulty of correctly classifying each voxel. Lower values imply higher classification difficulty. (e) depicts the final weights assigned to each voxel, where higher weights focus on rare, misclassified regions while down weighting well-classified background areas. This weighting strategy reduces the influence of abundant, correctly predicted voxels and increases the focus on difficult, misclassified lesions, thereby enhancing the model’s focus on challenging samples.
2.7 Model Evaluation
We assessed the performance of the loss functions on the overall test data. We made predictions on the PET/CT whole-body images using the sliding-window technique [29] with a window of dimension (128, 128, 128) for all networks. The test set predictions were resampled to the coordinates of the original ground truth masks to compute the evaluation metrics. In addition to assessing performance on the overall test set, we performed evaluations based on two different lesion scenarios, single and multiple lesion scenarios. First, we categorized the set of images ($I$) into two subsets: images with a single-lesion ($S$) and images with multiple-lesion ($M$). For an image $x$ with $n(G_{l})$ ground truth lesions, we defined these groups as follows:
$S=\{x\in I\mid n(G_{l})=1\}$
$M=\{x\in I\mid n(G_{l})\geq 2\}$
Then, for each scenario, as well as on the overall test set, we performed patient-level and lesion-level assessments based on the methodologies below.
2.7.1 Patient-Level Analysis
Segmentation Metrics:
We evaluated the segmentation performance of the loss functions using the DSC. We report the mean DSC together with the standard deviation and the median DSC with the inter-quartile ranges. Let $G$ and $P$ represent the ground truth and predicted masks of an image at the patient level. The patient-level DSC is defined as:
$\text{DSC}=\frac{2|G\cap P|}{|G|+|P|}$ (13)
Detection Metrics:
We defined true positive (TP), false positive (FP), and false negative (FN) detections at the patient level, reporting their mean values across the test set and computed F1 scores per patient. A detection is considered as TP if there is a matched pair of $G_{l}$ and $P_{l}$ such that $G_{l,\text{SUVmax}}\cap P_{l}\neq 0$, where $G_{l,\text{SUVmax}}$ is the voxel in $G_{l}$ with the maximum standardized uptake value (SUVmax) [30]. For a prediction $P_{l}$, for which there is no corresponding overlap with a ground truth lesion $G_{l}$, the detection is designated as an FP. Similarly, a false negative occurs when $G_{l,\text{SUVmax}}\cap P_{l}=0$. We provide a visual illustration of the definition of the detection metrics in Figure 2. The F1 score is thus computed as:
[…]
Figure 2: Illustration for defining a true positive detection based on an overlap with the voxel containing the maximum standardized uptake value (SUVmax) in the ground truth lesion. G is the set of ground truth lesions and P is the set of predicted lesions.
Clinical Metrics:
For each scenario, we analyzed the performance of the loss functions across different groupings of molecular tumor volume. Specifically, we evaluated the performance of the loss functions on DSC on different thresholds of total molecular tumor volume (TMTV). We computed thresholds $(t)$ based on the inter-quartile range from 0 to the 85th percentile of the ground truth TMTV values and calculated median DSC for all lesions $(l)$ where volume $(v_{l}>t)$. Additionally, for the multiple-lesion scenario only, we assessed the loss function performance based on the spatial extent of lesion distribution, $D_{\text{max}}$, which is measured as the maximum distance between any pair of foreground voxels in the image.
For a given ground truth mask, let $\mathbf{v}_{i}$ denote foreground voxels, where $i=1,2,\ldots,N$ and $N$ is the total number of lesion voxels. We calculated the Euclidean distance between every pair of lesion voxels, $\mathbf{v}_{i}$ and $\mathbf{v}_{j}$, accounting for voxel spacing in each dimension. For voxel coordinates $(x_{i},y_{i},z_{i})$ and $(x_{j},y_{j},z_{j})$ with spacing $(s_{x},s_{y},s_{z})$, the distance $d_{ij}$ is given by equation (15) below:
$d_{ij}=\sqrt{\left(s_{x}(x_{i}-x_{j})\right)^{2}+\left(s_{y}(y_{i}-y_{j})\right)^{2}+\left(s_{z}(z_{i}-z_{j})\right)^{2}}$ (15)
The lesion dissemination, $D_{\text{max}}$, is calculated as the maximum distance among all calculated distances; $D_{\text{max}}=\max_{i,j}\,d_{ij}$. We categorized the calculated distances into inter-quartile ranges (IQR) indicating the first, second, third and fourth IQR and analyzed the performance of the loss functions on each group.
2.7.2 Lesion-Level Analysis
Segmentation Metrics:
For lesion-level analysis, we defined the DSC based on the overlap between each predicted lesion mask and its corresponding ground truth mask. We identified which predicted lesion voxels corresponded to the ground truth voxels, especially in the scenario of multiple lesions in an image, based on a voxel-wise matching strategy. Specifically, lesions were segmented as individual connected components in both the ground truth and predicted masks using 18-connectivity to ensure spatial continuity within each identified lesion. This produced unique integer labels for each lesion. Each connected component in the predicted mask was then assessed for spatial overlap with each connected component in the ground truth mask.
For each pair of ground truth and predicted lesions, a match was defined if there was any voxel overlap between the components (i.e., if any voxels in a predicted lesion occupied the same spatial locations as those in a ground truth lesion). This was computed by checking if any intersecting voxels existed between the two labeled components. When an overlap was identified, the corresponding pair of ground truth and predicted lesion labels was recorded as a matched lesion pair. These matches were used to compute lesion-wise metrics, including lesion-level DSC, by comparing the voxel distributions in each matched pair of lesions. We defined the lesion-wise DSC as:
$\text{DSC}_{l}=\frac{2|G_{l}\cap P_{l}|}{|G_{l}|+|P_{l}|}$ (16)
where $G_{l}$ and $P_{l}$ denote the ground truth and predicted masks of a specific lesion. For a given ground truth lesion with no match, $|G_{l}\cap P_{l}|=0$ yielding a DSC of 00.
Clinical Metrics:
At the lesion level, analysis was performed only on the molecular tumor volumes of the lesions. Similar to the threshold analysis performed at the patient level, we calculated thresholds $(t)$ based on the inter-quartile range from 0 to the 85th percentile of the ground truth MTV values (molecular volume of individual lesions) and calculated median DSC for all lesions $(l)$ where volume $(v_{l}>t)$. For each scenario, we analyzed the performance of the loss functions across different groupings of MTV.
3 Results
3.1 Segmentation performance across different networks
In this section, we present results for the segmentation performance of the three loss functions, Dice Loss (DL), Dice Focal Loss (DFL), and L1-weighted Dice Focal Loss (L1DFL) using the Attention U-Net and SegResNet architectures (Table 1). We report the average and median DSCs at the patient level, with mean values of true positive, false positive, and false negative counts on the test set and the F1 score performance. A one-tailed paired Wilcoxon signed-rank test was performed to evaluate the significance of the performance of L1DFL from those of DL and DFL separately, with corresponding p-values reported.
Network Loss Function DSC Mean Rates F1 scores P values
Mean Median TP FP FN
Attention U-Net DL 0.54 $\pm$ 0.24 0.58 [0.44, 0.72] 0.81 $\pm$ 0.35 2.06 $\pm$ 1.89 0.19 $\pm$ 0.35 0.50 $\pm$ 0.28 $<0.01$*
DFL 0.51 $\pm$ 0.26 0.54 [0.33, 0.73] 0.77 $\pm$ 0.36 2.65 $\pm$ 2.60 0.23 $\pm$ 0.36 0.44 $\pm$ 0.27 $<0.01$*
L1DFL 0.58 $\pm$ 0.27 0.66 [0.51, 0.77] 0.77 $\pm$ 0.37 0.42 $\pm$ 0.65 0.23 $\pm$ 0.37 0.69 $\pm$ 0.34
SegResNet DL 0.52 $\pm$ 0.30 0.60 [0.29, 0.76] 0.73 $\pm$ 0.39 0.73 $\pm$ 1.10 0.27 $\pm$ 0.39 0.62 $\pm$ 0.35 0.24
DFL 0.52 $\pm$ 0.25 0.59 [0.41, 0.71] 0.81 $\pm$ 0.34 2.05 $\pm$ 2.00 0.19 $\pm$ 0.34 0.49 $\pm$ 0.27 0.015*
L1DFL 0.57 $\pm$ 0.29 0.68 [0.47, 0.78] 0.76 $\pm$ 0.37 0.52 $\pm$ 0.76 0.24 $\pm$ 0.37 0.66 $\pm$ 0.34
Table 1: Comparison of the loss functions on the test set based on mean and median patient-level DSC. The mean DSC, TP, FP, FN and F1 scores are presented with standard deviations and the median DSC with interquartile ranges. A detection is considered a true positive if the predicted mask overlaps with the voxel containing the SUVmax value in the ground truth. The P values reflect a one-tailed paired Wilcoxon signed-rank test comparing the median difference in DSC between L1DFL and DL and DFL, with * indicating statistical significance at the $\alpha=0.05$ level.
The median DSCs are higher than the average DSCs, with SegResNet slightly outperforming Attention U-Net in DSC. However, Attention U-Net generally got better true positive rates except in the case of DFL, where SegResNet performed better with a $5\%$ increment. In both architectures, Dice Loss always performed better than Dice Focal Loss in minimizing false positives. However, both loss functions showed varied performances on the different networks regarding true positive rates. Specifically, while the Dice Loss surpassed the other loss functions with a true positive rate of 0.81 $\pm$ 0.35 on the Attention U-Net, Dice Focal Loss also achieved a similar performance but on the SegResNet architecture. Yet, this higher sensitivity of Dice Loss on Attention U-Net came at the cost of more false positives. Compared to a false positive rate of 0.73 $\pm$ 1.10 on SegResNet, the Dice Loss had a rate of 2.06 $\pm$ 1.89 on Attention U-Net, about three times the value on the former network. This imbalance in true and false positives on Attention U-Net led to a lower F1 score for Dice Loss despite its higher sensitivity.
Conversely, L1DFL achieved the best balance of true positive detections and minimum false positives. For example, on Attention U-Net, where Dice Loss and Dice Focal Loss had mean false positive rates above 2.0, the L1DFL minimized the false detection rates. Although its FP rate slightly increased with SegResNet, its value was still low compared to DL and DFL. By lowering false positive rates while maintaining sensitivity to lesions, L1DFL yielded superior mean and median DSCs. Specifically, on Attention U-Net, L1DFL reached a median DSC of 0.66, outperforming DFL and DL by $22\%$ and $13\%$, respectively. Moreover, the high interquartile range of L1DFL indicates that its DSC values fell within a narrow range of high values.
Although the one-tailed Wilcoxon signed-rank test comparing DL and L1DFL did not provide enough evidence to reject the null hypothesis of no significant difference between the two on the SegResNet architecture, statistically significant differences were obtained on the Attention U-Net. Also, comparing DFL versus L1DFL resulted in statistically significant differences in favor of L1DFL across both networks.
Figure 3: Segmentation performance of the loss functions across two networks: Attention U-Net (top row) and SegResNet (bottom row). The first column shows the ground truth segmentation (GT) and the remaining columns represent the results for the loss functions: Dice Loss (DL), Dice Focal Loss (DFL), and L1-weighted Dice Focal Loss (L1DFL). The green-colored region denotes the ground truth annotation of the lesion, while the red-colored regions represent the segmentations produced by the models. The bounded area highlights the lesion in the ground truth image alongside the corresponding correct segmentation by the models. Red regions outside the bounded area indicate false positive segmentations.
Figure 3 shows a sample segmentation performance of the loss functions on a test PET/CT image. The first-row segmentations are from Attention U-Net, while the second row is from SegResNet, with columns having different loss functions. The bounded region in the figure depicts correct segmentation according to the ground truth highlighting the differences in performance between the loss functions and the network architectures.
In this example, Attention U-Net yielded better segmentation of the lesion with higher DSC than SegResNet. Yet, for both networks, the Dice and the Dice Focal Loss functions produced instances of false positives further impacting the overall DSC. L1DFL, however, remained robust, providing a consistent performance.
3.2 Performance in single and multiple lesion scenarios
We evaluated the performance of the loss functions in two different lesion scenarios: multiple lesions in one PET/CT volumetric image and only one lesion in the volume constituting the single lesion scenario. Figure 4 shows that the loss functions performed differently under these scenarios; a wider spread of DSC values is observed in the single lesion scenario than in the multiple scenario for both networks. Also, the DSCs beyond the 75th percentile for each loss function in the single lesion scenario were higher than those in the multiple lesion scenario, with more cases achieving DSCs greater than 0.8.
Despite this observation, each loss function with Attention U-Net had higher median DSCs in the multiple-lesion scenario than in the single-lesion scenario (Figure 4a). Besides, in the single lesion case, all the loss functions have more lower outliers, reflecting occasional underperformance. On the other hand, for all the loss functions, there are fewer outliers in the multiple-lesion scenario. While similar observations are made on the SegResNet architecture, the median DSC of L1DFL in the single lesion scenario remained slightly higher than the median DSC in the multiple lesion case (Figure 4b).
Figure 4: Box plots overlaid with swarm plots illustrating the distribution of Dice Similarity Coefficient (DSC) performance for the three loss functions across two lesion scenarios: M (lesion count $>2$) and S (lesion count $=1$). Each box plot’s top and bottom edges represent the interquartile range (IQR), while the horizontal lines within the boxes indicate the median DSC values. Whisker lengths are set to 1.5 times the IQR. (a) Performance on Attention U-Net architecture. (b) Performance on SegResNet. The number of cases in the single lesion scenario was N = 34 and for the multiple lesion scenario, the number of cases was N = 23.
Dice Loss generally had higher median DSC for both architectures when compared with DFL, although there were variations in performance between lesion scenarios and architectures. In a single lesion scenario using the Attention U-Net architecture, for example, DL demonstrated more consistent performance as evidenced by an IQR of 0.33 compared to DFL with an IQR of 0.45. However, on SegResNet, DL’s performance became highly incoherent, with a wider IQR of 0.72 compared to the DFL’s narrower IQR of 0.37.
In the case of multiple lesions, while both loss functions showed similar IQRs of 0.26 with SegResNet, on Attention U-Net, however, the DSC values from DL are more spread out with lower values in the lower quartile range compared to DFL and L1DFL. Thus, Dice Loss may not be consistent in some scenarios and architectures, particularly where the cases are more complex.
In contrast, L1DFL outperformed DL and DFL in both lesion scenarios on the Attention U-Net. Its median DSCs on the multiple and single lesions were 0.67 and 0.65, respectively. These values are at least $11\%$ higher than the median DSCs of DL in both cases. Similarly, on SegResNet in the single lesion case, L1DFL achieved a median DSC of 0.69, while DL and DFL had 0.56 and 0.55, respectively. However, in the multiple lesion scenario, Dice Loss slightly outperformed L1DFL by $4\%$, although they both had similar IQR.
3.3 Performance based on lesion molecular volume
For each lesion scenario, we evaluated the performance of the loss functions across different TMTV thresholds at both the patient and lesion levels (Figure 5). Higher DSC values were generally observed when assessing the performance at the lesion level while the patient-level DSC values tended to be slightly lower due to false-positive segmentations. Notably, the Dice and Dice Focal Losses were more susceptible to false positives, resulting in a reduction of their median DSC by at least $10\%$ in the patient-level assessment.
Figure 5: Patient and lesion-level median Dice Similarity Coefficient (DSC) performance of the different loss functions as a function of various ground truth tumor molecular volume thresholds. Each column shows the plots for the different lesion scenarios: single (S) and multiple (M). At both the patient and lesion levels, the line plots represent the trend of each loss function’s performance across different total molecular tumor volume (TMTV) and molecular tumor volume (MTV) thresholds, respectively, with thresholds $(t)$ based on the interquartile range from 0 to the 85th percentile. Median DSC is calculated for all lesions $(l)$ where volume $(v_{l}>t)$, and each color indicates a different loss function. The first row of plots (a) shows the performance of the loss functions with Attention U-Net. The second row (b) shows the performance of the loss functions with SegResNet. The third row (c) displays area plots of median tumor lesion activity (TLA) values according to the TMTV and MTV thresholds, where median TLA values are calculated for lesions with $(v_{l}>t)$.
[…]
The Dice Loss performed better on SegResNet than on Attention U-Net, outperforming Dice Focal Loss at the patient level in both scenarios due to Dice Focal Loss’ higher false positive rate. However, both loss functions showed similar trends at the lesion level. The Dice Focal Loss on the other hand yielded more accurate segmentations with the Attention U-Net outperforming the Dice Loss at both scenarios. Nevertheless L1DFL maintained stable performance across both architectures and lesion scenarios. This consistency suggests the effectiveness of L1DFL’s dynamic weighting strategy, making it more robust to false positives at the patient level and achieving better overlap with the ground truth at the lesion level.
3.4 Performance based on lesion dissemination
To quantify the spatial distribution of lesions, we calculated $D_{\text{max}}$, the maximum Euclidean distance between any two lesions in the image.
Figure 6: Plots illustrating the Dice Similarity Coefficients (DSC) across four interquartile ranges of $D_{\text{max}}$. The first row (a) shows the performance with Attention U-Net. The second row (b) shows the performance with SegResNet. The groups G0 (0-9 cm), G1 (9-11cm), G2 (11-14cm), and G3 (14-60cm) correspond to the interquartile ranges of $0-25\%$, $25-50\%$, $50-75\%$, and $75-100\%$, respectively. In the box plots, horizontal lines indicate the median, and whisker lengths are set to 1.5 times the interquartile range (IQR), with outliers represented as black dots. Accompanying bar plots (c) display the median total lesion activity (TLA) values for cases within each $D_{\text{max}}$ group. The different loss functions are color-coded for clarity. The total number of cases in each group was N = 6.
For the performance assessment of the loss functions based on the $D_{\text{max}}$, we categorized the distances into the following groups: 0-9 cm (G0), 9-11 cm (G1), 11-14 cm (G2), and 14-60 cm (G3), representing the first, second, third and fourth quartile ranges, respectively. Patient-level analysis of DSC showed consistent improvement across these ranges (Figure 6). The figure also shows differences in performance across the different architectures. Generally, the Dice Loss performed better on SegResNet than on Attention U-Net but the Dice Focal Loss performed better on the latter architecture. The L1DFL on the other hand, yielded more consistent performance across the two networks, yet had a wider IQR on SegResNet than on Attention U-Net.
On Attention U-Net, initially, median DSC values for all loss functions ranged from 0.4 to 0.6 (G0), increasing in the third quartile to 0.6-0.8 (G2). The results with SegResNet also show a similar trend; there is a general increase in DSC values up to the third quartile. For both networks, in the last quartile (G3), representing a more expansive lesion spread, all loss functions showed a drop in their performance metrics. Specifically, on Attention U-Net, the Dice Focal Loss performed the worst in this group, and L1DFL and Dice Loss presented similar performances, with a median of 0.62. Yet, the Dice Focal Loss had a smaller interquartile range, reflecting less spread across DSCs in that category. The median DSC of L1DFL dropped from 0.76 to 0.62 by $18.4\%$, and that of Dice Focal Loss dropped from 0.65 to 0.58 by $10.3\%$. Dice Loss also slightly declined in performance with a broader spread in IQR of its DSCs. On SegResNet, the Dice Loss had the worst performance in this group with L1DFL achieving the best segmentation accuracy.
The performance of the loss functions in the lower $D_{\text{max}}$ ranges (G0 - G2) which characterize less lesion spread in the body further highlight network-specific variations. While the Dice Loss achieved similar high median DSC as L1DFL in these ranges with the SegResNet, it had the worst performance with Attention U-Net. Similarly, the SegResNet model trained with the Dice Focal Loss had a decline in obtaining accurate segmentations while pairing this loss function with Attention U-Net yielded competitive results. L1DFL remained robust to these architecture-specific variations consistently outperforming the baseline models. Figure 6 also depicts a higher median total lesion activity (TLA) across the different groups, indicating an increased lesion size or uptake value with greater lesion spread. Although model performance generally improves with higher lesion activity, the slight decrease in median DSC in the highest quartile of $D_{\text{max}}$, despite high TLA, suggests that other factors might contribute to the segmentation performance such as lesion site and count.
4 Discussion
In this work, we proposed and evaluated a novel loss function L1-weighted Dice Focal Loss, in segmenting metastatic prostate cancer lesions based on different lesion scenarios characterized by the number of lesions in an image. Specifically, we defined two scenarios: single lesion scenario which include whole body PET/CT images with only a single lesion present in the ground truth mask and multiple lesion scenario being images with more than one lesion. We compared the performance of our proposed loss function with the Dice and Dice Focal Losses. Moreover, we report how the size of a lesion and its spatial distribution influence the segmentation result. The dataset used in this study included tumors which had resurfaced after treatment with relatively low molecular volumes and standardized uptake values (SUV) thus introducing unique challenges for correct segmentations.
Our findings showed that the performance of the loss functions is not uniform across different architectures. For instance, the results show that the Dice Loss generated more false positives on Attention U-Net compared to SegResNet, despite showing higher sensitivity on the former. Similarly, the Dice Focal Loss exhibited lower precision on the Attention U-Net architecture due to increased false positives. By evaluating performance at both the patient and lesion lesions, the effect of false positive rates was highlighted. There was a more than $10\%$ decrease in the median DSC values at the lesion level for both loss functions when assessments were made at the patient level. L1DFL, in contrast, was more robust to these false detections.
[…]
Furthermore, comparing the results obtained with SegResNet on the single lesion scenario to the multiple lesion scenario points to the influence of additional factors to segmentation accuracy other than tumor volume and activity. Due to the summation of the TLA values of individual lesions in the multiple lesion scenario, higher activity values are obtained than in the single lesion case yet, lower DSC values are observed. This observation could suggest the introduction of further complexities as lesion count increases. In fact, with more lesions, different anatomical sites may be involved, and the effect of other organs with high uptake values like the bladder could be pronounced [31]. Besides, the evaluation based on lesion spread, $D_{\text{max}}$, highlights the decrease in the models’ performance as distribution across the body becomes more expansive.
The Dice Loss is based on overlap measurement and it inherently balances sensitivity with specificity, making it more robust in scenarios with significant data imbalance such as in lesion segmentation tasks [16]. Therefore, distribution-based losses are often combined with the Dice Loss or its variants to improve their robustness. However, without adapting its focus to classification difficulty, the Dice Loss fails to adjust dynamically toward heterogeneous lesion characteristics. Despite the unique challenges posed by the tumor heterogeneity, the L1-weighted Dice Focal Loss generally yielded better segmentation performance, achieving better balance in false detection rates. This loss function comprises a dynamic weighting strategy that adapts the model’s focus based on sample difficulty. For easier-to-classify examples (characterized by predicted probabilities $P>0.5$), L1DFL applies a reduced penalty while penalizing more challenging examples ($P<0.5$) more heavily (Figure 7).
Figure 7: Plot of loss versus probability for samples with the ground truth label as 1. The curves represent the behavior of each loss function, highlighting how they respond to varying probabilities of correctly classifying examples. The plot illustrates the differences in how each loss function handles both easy-to-classify and hard-to-classify cases based on their predicted probabilities.
Similarly, the Dice Focal Loss increases the penalty on samples with a lower probability of correctly classifying the ground truth label, P(G), encouraging the model to focus more on ”hard” samples, but it applies a much steeper penalty. Thus, it tends to lose the Dice Loss component, where positive and negative examples remain equally weighted, potentially leading to more false positive rates due to misclassification of background areas as lesions [18]. On the other hand, the weighting strategy of L1DFL allows for a more balanced approach to penalizing misclassification, consequently reducing false positive rates. Our results indicate that L1DFL yielded the highest F1 and DSCs in this study, underlining its more robust performance across various lesion scenarios and levels of lesion dissemination. This adaptive weighting strategy shows promise in improving segmentation performance in complex metastatic imaging cases.
5 Conclusion
We evaluated the performance of a custom loss function, L1-weighted Dice Focal Loss (L1DFL), in the detection and segmentation of metastatic prostate cancer lesions in PSMA PET/CT scans across single and multiple lesion scenarios, characterized by lesion count. L1DFL leverages a dynamic weighting scheme that is based on the L1 norms of the predicted probabilities and ground truth labels to modulate the Dice Loss and then combined with the Focal Loss to emphasize challenging pixel classifications. We compared L1DFL’s performance against Dice Loss and Dice Focal Loss using two 3D CNN architectures, Attention U-Net, and SegResNet. While Dice Loss demonstrated sensitivity for small lesions, it usually suffered from low precision. Dice Focal Loss showed higher false positive rates and variable performance across the different networks. On the other hand, L1DFL consistently achieved a balanced segmentation performance across the lesion scenarios and architectures, achieving superior F1 scores and Dice Similarity Coefficients. Future work will be directed at exploring the generalizability of L1DFL to other medical imaging datasets and tasks.
6 Acknowledgments
This work was supported by the Canadian Institutes of Health Research (CIHR) Project Grants PJT-162216 and PJT-173231. We also acknowledge computational resources and services provided by Microsoft AI for Health.

### v2

@@
 Abstract
-This study proposes a new loss function for deep neural networks, L1-weighted Dice Focal Loss (L1DFL), that leverages L1 norms for adaptive weighting of voxels based on their classification difficulty, towards automated detection and segmentation of metastatic prostate cancer lesions in PET/CT scans.
-We obtained 380 prostate-specific membrane antigen (PSMA) [18F]DCFPyL PET/CT scans of patients diagnosed with biochemical recurrence metastatic prostate cancer.
-We trained two deep neural networks with 3D convolutional neural network backbone — Attention U-Net and SegResNet — and concatenated the PET and CT volumes channel-wise as input.
-The performance of our custom loss function was evaluated against the Dice and Dice Focal Loss functions.
-For clinical significance, we considered a detected region of interest (ROI) as a true positive if at least the voxel with the maximum standardized uptake value falls within the ROI.
-We assessed the models’ performance based on the number of lesions in an image, tumor volume, activity and extent of lesion spread.
-The L1DFL outperformed the comparative loss functions by at least $13\%$ on the test set.
-In addition, the F1 scores of the Dice Loss and the Dice Focal Loss were lower than that of L1DFL by at least $6\%$ and $34\%$, respectively.
-The Dice Focal Loss yielded more false positives, whereas the Dice Loss was more sensitive to smaller volumes and struggled to segment larger lesions accurately.
-They also exhibited network-specific variations and yielded declines in segmentation accuracy with increased tumor spread.
-Our results demonstrate the potential of L1DFL to yield robust segmentation of metastatic prostate cancer lesions in PSMA PET/CT images.
-The results further highlight potential complexities arising from the variations in lesion characteristics that may influence automated prostate cancer tumor detection and segmentation.
-The code is publicly available at: https://github.com/ObedDzik/pca_segment.git.
+Background and Objective: Accurate automated detection of recurrent prostate cancer in PSMA PET/CT scans is challenging due to heterogeneous lesion size, activity, anatomical location, and intra- and inter-class imbalances.
+Conventional deep learning loss functions often produce suboptimal optimization, as gradients are dominated by easy background voxels or extreme outliers.
+To address this, we propose L1-weighted Dice Focal Loss (L1DFL), which harmonizes gradient magnitudes across voxels using L1 norms to adaptively weight samples based on classification difficulty, resulting in well-calibrated predictions with a bimodal separation between correct and incorrect predictions.
+Methods: We trained three 3D convolutional networks (Attention U-Net, SegResNet, U-Net) and a transformer-based UNETR model on 380 PSMA PET/CT scans.
+PET and CT volumes were concatenated as input.
+A finetuned medical foundation model (SAM-Med3D) was also evaluated.
+L1DFL was compared against Dice Loss (DL) and Dice Focal Loss (DFL) using metrics including Dice Similarity Coefficient (DSC), F1 score, and lesion-wise detection performance, with statistical significance assessed via Wilcoxon signed-rank tests with Bonferroni correction ($\alpha=0.003$).
+Results: Across architectures, L1DFL consistently outperformed DL and DFL, achieving at least a $4\%$ improvement in DSC.
+F1 scores were higher by $\geq 6\%$ and $\geq 26\%$ compared to DL and DFL, respectively.
+While DFL produced more false positives and DL struggled with larger lesions, L1DFL achieved balanced detection, minimizing false detections while maintaining high true positive rates.
+The gradient harmonization mechanism ensured robustness across varying lesion sizes, volumes, and spread.
+Conclusion: L1DFL improves segmentation of recurrent prostate cancer lesions in PSMA PET/CT by harmonizing gradients.
+It produces well-calibrated predictions with clear separation of correct and incorrect cases, and balanced detection across lesion sizes while maintaining consistent performance across multiple architectures.
 keywords:
@@
 […]
-[a]organization=School of Biomedical Engineering,addressline=University of British Columbia, city=Vancouver, country=Canada
+[a]organization=University of British Columbia,city=Vancouver, country=Canada
 […]
-[b]organization=Department of Integrative Oncology,addressline=BC Cancer Research Institute, city=Vancouver, country=Canada
+[b]organization=BC Cancer Research Institute,city=Vancouver, country=Canada
 […]
 [c]organization=Vector Institute,city=Toronto, country=Canada
+{highlights}
+Besides uptake, tumor characteristics such as volume and dissemination limit automated detection performance.
+The proposed loss function demonstrates strong calibration, producing a clear separation between correct and incorrect predictions.
+It dynamically distributes gradient magnitudes across samples, avoiding overemphasis on outliers or overly easy examples.
+It outperforms the Dice and Dice Focal Losses by at least $4\%$ based on Dice Score.
+It outperforms Dice Loss and Dice Focal Losses by $6\%$ and $26\%$, respectively, on F1 score.
+1 Introduction
+Semantic segmentation assigns a label to every pixel or voxel in an image [9], requiring both local feature extraction and broader contextual understanding [5], since spatial coherence and boundary smoothness are critical [34].
+In medical imaging, segmentation accuracy is particularly critical, as false detections or missed lesions can significantly affect diagnosis, staging, and treatment planning [28].
+These difficulties are compounded by severe class imbalance, where pathological voxels constitute only a small fraction of the image volume.
+In whole-body oncological PET, lesion voxels are heavily outnumbered by background, and lesions vary widely in size, shape, contrast, and tracer uptake [15].
+Metastatic prostate cancer (mPCa) amplifies these challenges further, with lesions distributed across multiple anatomical sites [10].
+PSMA PET imaging adds another layer of complexity, as PSMA expression appears in normal organs and inflammatory processes as well as malignant tissue [19].
+AI-based detection systems must therefore balance sensitivity to small or low-uptake lesions against specificity to avoid false positives from physiological uptake.
+Most automated approaches struggle with this trade-off.
+Techniques such as standardized uptake value (SUV) thresholding are commonly employed post hoc to suppress false positives by classifying voxels above a predefined threshold as foreground [11].
+While effective in reducing spurious detections, such strategies often sacrifice sensitivity for lesions with low tracer avidity or lesions in patients with atypical uptake patterns [23].
+More sophisticated methods have attacked the problem at the input or representation level: semi-automatic pipelines with bone masking and SUV thresholding [12], nnU-Net-based whole-body frameworks [16, 17], anatomical organ priors [23], 2D diffusion models on MIP projections [37], and hybrid MIP-based frameworks [7].
+Feature-clipping normalization [3] and 2.5D pipelines [44] have also been explored.
+While each has yielded incremental gains, false positives and missed small or ambiguous lesions remain persistent failure modes, and critically, none directly addresses how voxel-level errors are weighted during optimization.
+In PET lesion segmentation, not all voxels contribute equally to learning.
+The majority of background voxels are trivially classified, while a small subset of foreground and boundary voxels is consistently misclassified or uncertain.
+Standard losses do not distinguish between these difficulty levels.
+Region-based losses such as Dice Loss (DL) handle inter-class imbalance through overlap-based formulation [8, 42] but treat all voxels within a class equally [43].
+Distribution-based losses like Focal Loss [24] and TopK loss [39] address difficulty-based weighting but rely on fixed hyperparameters that limit adaptability to mini-batch distributions [26], and improper weighting can compromise sensitivity-specificity balance [41].
+Compound losses such as Dice Focal Loss [45] and Combo Loss [36] leverage complementary strengths but inherit the same static weighting limitations.
+A more principled solution is gradient harmonization [22], which adaptively weights samples based on gradient density rather than fixed hyperparameters, preventing easy examples from overwhelming the contribution of hard minority cases.
+Liu et al.
+[26] extended this to medical image segmentation via knee MRI, demonstrating its value for handling intra-class imbalance.
+However, this approach has not yet been explored for PET lesion segmentation, where the gradient density problem is especially acute.
+We therefore propose L1-weighted Dice Focal Loss (L1DFL), which integrates gradient harmonization into the Dice Focal Loss framework.
+Unlike prior approaches that bin samples by computed gradient magnitudes [22, 26], L1DFL uses the L1 norm between predicted probabilities and ground truth labels as a canonical proxy for voxel-level classification difficulty [21, 33, 25].
+Voxels in high-density difficulty regions are down-weighted while those in low-density regions are up-weighted, distributing gradient contributions uniformly across the difficulty spectrum and avoiding domination by any single difficulty level during optimization.
+The main contributions of this work are: (i) a novel loss function, L1DFL, incorporating gradient harmonization; (ii) evaluation of L1DFL against DL and DFL on false positive and false negative rates; (iii) comparison across single-lesion and multiple-lesion scenarios; (iv) assessment on clinical metrics including molecular tumor volume and lesion spread; and (v) benchmarking across five models: Attention U-Net [32], SegResNet [30], U-Net [6], UNETR [14], and SAM-Med3D [38].
+2 Materials and Methods
+2.1 Dataset and Ground Truth Annotation
+We analyzed 380 [18F]DCFPyL PET/CT scans of patients diagnosed with biochemical recurrence of prostate cancer as part of an ongoing clinical trial (NCT02899312) with informed written consent obtained from the participants.
+The image data collection was approved by the University of British Columbia – BC Cancer Ethics Board.
+The inclusion criteria for the patients were: (1) histologically proven prostate cancer with biochemical recurrence after initial curative therapy with radical prostatectomy, with a PSA >> 0.4 ng/mL and an additional PSA measurement confirming an increase; and (2) histologically proven prostate cancer with biochemical recurrence after initial curative radiotherapy, with a PSA >> 2 ng/mL after therapy, with five or fewer lesions identified on [18F]DCFPyL PET/CT [13].
+Patients received an average of 350 MBq of [18F]DCFPyL following a four-hour fast, with dose adjustments based on body weight.
+PET/CT imaging was performed 120 minutes post-injection using GE Discovery 600 or 690 scanners (GE Healthcare, USA).
+A non-contrast CT scan was acquired for attenuation correction and anatomical localization (120 kV, automated mA modulation), followed by a whole-body PET acquisition lasting 2-4 minutes per bed position.
+PET images were reconstructed using ordered subset expectation maximization with point-spread function modeling, producing transaxial images with a matrix resolution of $192\times 192$ and an in-plane pixel spacing of $3.64~\text{mm}$ x $3.64~\text{mm}^{2}$ and a slice thickness of $3.27~\text{mm}$ [13].
 […]
-[d]organization=Department of Physics & Astronomy,addressline=University of British Columbia, city=Vancouver, country=Canada
-[…]
-[e]organization=Department of Radiology,addressline=University of British Columbia, city=Vancouver, country=Canada
-[…]
-[f]organization=Department of Electrical and Computer Engineering,addressline=University of British Columbia, city=Vancouver, country=Canada
-1 Introduction
-Semantic segmentation involves mapping every pixel of an image to a class, whether an object or a background [1].
-It requires both a local and contextual understanding of the image [2], making it particularly difficult since neighboring pixels are highly related; thus, information on the spatial context and the boundary’s smoothness is essential in a segmentation model [3].
-Particularly in medical image analysis, higher accuracy levels are required for tumor segmentation since false detections may significantly impact patient management [4].
-However, medical images are fraught with an imbalance in the proportion of voxels belonging to the regions of interest (ROIs) compared to the background making a balance in precision and sensitivity in automated segmentation a challenging task.
-Whole body positron emission tomography (PET) scans of cancer patients are an example of such medical images where voxels corresponding to the foreground class are overwhelmingly outnumbered by background voxels.
-These images are characterized by high variability in lesion size, shape, and intensity, as well as significant heterogeneity in tracer uptake [5].
-In the case of metastatic prostate cancer (mPCa) images, lesions can range from small, low-contrast regions barely distinguishable from surrounding tissues to large, high-uptake areas with well-defined boundaries, and may occupy different anatomical sites [6].
-In localizing these lesions, the prostate-specific membrane antigen (PSMA) is often targeted due to its high expression by tumor cells.
-However, PSMA expression is not exclusive to prostate cancer lesions; it is also found in other tissues, including salivary glands, kidneys, and the small intestine, as well as in certain non-prostatic malignancies and inflammatory processes [7].
-This broad expression can lead to radiotracer uptake in non-cancerous or non-prostatic regions challenging the specificity of computer-aided detection algorithms.
-Thus, the challenge in artificial intelligence (AI) based lesion detection and segmentation in PET scans of metastatic prostate cancer lesions is avoiding false-positive segmentation of anatomical structures and other artifacts while maintaining accurate lesion detection.
-With most of the automated approaches, improved sensitivity of mPCa lesions often comes at the cost of specificity, yet both extremes pose potential risks.
-One commonly used strategy for limiting false positive rate is standardized uptake value (SUV) thresholding, which involves defining all voxels with SUV greater than some value, typically considering the background uptake of the liver to be foreground and all other voxels as background [8].
-This method may be included as part of the automated segmentation process or as a post-processing step and has been shown to reduce false positive counts [9].
-Yet, its major drawback is the potential of having a reduced sensitivity for lesions that exhibit lower avidity for the radiotracer, or in patients with extensive liver metastases [10].
-Loss function design is another strategy that has received much attention and has led to an improvement in the performance of AI-based segmentation methods.
-The widely used loss function is the Dice Loss.
-It is based on the Dice Similarity Coefficient (DSC) and penalizes pixel-wise discrepancies between the predicted maps and their corresponding ground truth for each class [11].
-A model trained using the Dice Loss objective maximizes DSC inherently incorporating a balance in sensitivity and specificity [12].
-Nevertheless, the Dice Loss is not well-suited for handling another type of imbalance caused by a difference in the difficulty in classifying pixels [13].
-Within a class, be it the minority foreground class or majority background class, some pixels might be easier to classify, with models achieving greater than $50\%$ probability of correctly classifying such pixels [13].
-However, others may be more challenging to predict, yielding lower probabilities [13] and are therefore designated hard to classify.
-Without a focusing strategy implemented, the Dice Loss can be easily overwhelmed by the easy-to-classify samples and may perform poorly on the harder ones [14].
-Besides, it exhibits an inconsistent performance when segmenting multiple lesions of varying sizes in an image [15].
-These limitations have prompted the development of alternative loss functions.
-Distribution-based loss functions like Focal and Cross-Entropy losses are frequently used alternatives due to their dynamic focusing ability.
-They are often combined with the Dice Loss and its variants to form compound loss [11] since the losses based on overlap measurements are more robust for medical segmentation tasks [16].
-With an additional ability to control the model’s focus on detecting ROIs, these combo losses generalize well [11, 17].
-The challenge, however, remains on the balance of sensitivity and specificity as these compound losses tend to lose the Dice Loss component, potentially leading to more false positive segmentation [18].
-Thus, various regularizations and additional constraints are required to improve model accuracy for a medical lesion segmentation task [19].
-This work proposes a new loss function, L1-weighted Dice Focal Loss (L1DFL), with a dynamic weighting strategy based on L1 norms of predicted probabilities and ground truth labels.
-Using L1 norms, the loss function considers the difficulty of each sample and the prevalence of similar samples in the dataset.
-In medical image segmentation tasks, datasets are often imbalanced (disproportionately higher number of background voxels than ROIs) and dominated by easy samples (voxels that are easily classified correctly); the L1 norms of these easy samples tend to approach zero.
-Our technique focuses on down-weighting both easy samples and those with high L1 norm counts, which are likely to represent the background.
-We demonstrate the efficiency of L1DFL by experiments on [18F]DCFPyL PET/CT scans of patients with biochemical recurrence metastatic prostate cancer, for which lesions have varied sizes, primarily small, with low uptake values.
-The main contributions of this work are as follows: (i) we introduce a novel loss function, L1-weighted Dice Focal Loss (L1DFL), incorporating a dynamic weighting strategy; (ii) we evaluate the performance of L1DFL, Dice Loss, and Dice Focal Loss functions in handling false positives and false negative rates; (iii) we compare the segmentation performance of the loss functions based on different lesion scenarios: single-lesion, and multiple lesions.
-These scenarios are based on the number of lesions in an image; (iv) we evaluate the performance of these loss functions on different clinical metrics like molecular tumor volumes and the extent of lesion spread in the body; and (v) we assess performance using two network architectures, Attention U-Net [20] and SegResNet [21].
-2 Materials and Methods
-2.1 Dataset
-We analyzed 380 prostate-specific membrane antigen (PSMA) [18F]DCFPyL PET/CT scans of patients diagnosed with biochemical recurrence metastatic prostate cancer as part of an ongoing clinical trial (NCT02899312) with informed written consent obtained from the subjects.
-The image data collection was approved by the University of British Columbia – BC Cancer Ethics Board.
-The inclusion criteria for the patients were: (1) histologically proven prostate cancer with biochemical recurrence after initial curative therapy with radical prostatectomy, with a PSA $>$ 0.4 ng/mL and an additional PSA measurement confirming an increase; and (2) histologically proven prostate cancer with biochemical recurrence after initial curative radiotherapy, with a PSA $>$ 2 ng/mL after therapy, with five or fewer lesions identified on [18F]DCFPyL PET/CT [22].
-Each patient received an average dose of 350 MBq [18F]DCFPyL after fasting for four hours, with dose adjustments based on body weight.
-Imaging was conducted 120 minutes after injection using a GE Discovery 600 or 690 scanner (GE Healthcare, USA).
-A non-contrast-enhanced CT scan was acquired for localization and attenuation correction (120 kV, automatic mA adjustment from 30–200 mA, and a noise index of 20).
-Following CT acquisition, a whole-body PET scan was conducted at 2–4 minutes per bed position, depending on the participant’s size, with reconstruction performed using an ordered subset expectation maximization algorithm and point-spread function modeling.
-Each transaxial PET image had a matrix resolution of 192 × 192 pixels, with a physical pixel size of $3.64\text{mm}^{2}$.
-An expert nuclear medicine physician performed manual segmentation on all active lesions [22].
-There were a total of 684 prostate cancer (PCa) lesions across the entire dataset with a mean active lesion volume of 6.68 $\pm$ 10.20 ml.
-The lesions varied in uptake with the average maximum standardized uptake value (SUVmax) and mean standardized uptake value (SUVmean) being 12.65$\pm$14.46 and 4.62 $\pm$ 3.88, respectively.
-2.2 Ground Truth Annotation
-The ground truth segmentations were performed by a board-certified nuclear medicine physician from the BC Cancer Research Institute.
-All lesions were annotated using the PET Edge tool, a semi-automated gradient-based segmentation tool, and contours refined using the 3D Brush tool.
-Both tools are available in the MIM workstation (MIM software, Ohio, USA).
-These segmentations served as the ground truth labels for training the models.
 2.2 Data Preprocessing and Augmentation
 […]
-2.4 Architectures
-In this work, we implemented and studied two convolutional neural network architectures: Attention U-Net [20] and SegResNet [21].
-These networks, built on top of a 3D-CNN backbone, comprise encoder and decoder blocks.
-The SegResNet uses residual blocks with skip connections for feature propagation and maintaining gradient stability during training [21].
-The encoder path contained blocks with the following number of layers: 1, 2, 2, and 4, allowing for increasing levels of feature extraction at different resolutions.
-The initial filter size was set to 16 channels, then doubled through each downsampling step to capture progressively higher-level features.
-Each convolutional layer employs a ReLU activation function, and Group Normalization [23].
-The decoder path, containing one layer in each block, gradually reconstructs the segmented output while leveraging skip connections to preserve spatial details from the encoder.
-The input and output channels were set to 2, accommodating multi-channel input images and dual-class segmented outputs.
-On the other hand, Attention U-Net extends the U-Net architecture to include attention gates along the encoder-decoder path, allowing the network to selectively pay attention to essential features [20].
-The encoder path consisted of five layers, with each layer increasing in channel depth from 16, 32, 64, 128, to 256 channels, and each layer downsampled by a stride of 2.
-These attention gates guide the network in emphasizing regions that contribute significantly to the segmentation task.
-Each convolutional layer employs a ReLU activation function, and Batch Normalization as in the original implementation [20].
-The output layer matched the input with two channels for multi-modal inputs and dual-class output, representing background and foreground classes.
-2.5 Model Training
+2.3 Implementation Details
+Architectures:
+We evaluated five volumetric segmentation architectures: SegResNet [30], Attention U-Net [32], 3D U-Net [6], UNETR [14], and SAM-Med3D [38].
+SegResNet, Attention U-Net, and 3D U-Net were implemented as convolutional encoder–decoder networks with multi-channel inputs and two-class outputs.
+SegResNet employed encoder depths of {1, 2, 2, 4} with an initial channel size of 16, doubling at each downsampling stage, and used Group Normalization and ReLU activations.
+Attention U-Net consisted of five encoder levels with channel sizes {16, 32, 64, 128, 256}, Batch Normalization, and attention gates as originally proposed.
+The 3D U-Net comprised six levels with channels increasing from 16 to 512, residual blocks with Batch Normalization, and a symmetric decoder with transposed convolutions.
+UNETR and SAM-Med3D represent transformer-based and foundation-model approaches.
+UNETR used a vision transformer encoder with hidden size 256, MLP dimension 1024, and four attention heads, convolutional patch embeddings, and a U-shaped convolutional decoder.
+SAM-Med3D is a fully three-dimensional extension of the Segment Anything Model designed for volumetric medical image segmentation.
+It follows the original SAM design, consisting of a 3D image encoder, a 3D prompt encoder, and a mask decoder, with all 2D operations replaced by 3D counterparts, including 3D convolutions, normalization layers, and positional encodings [38].
+We used the publicly released SAM-Med3D pretrained weights and fine-tuned the model using our proposed and comparative loss functions.
+Model Training:
 We divided the dataset into training, validation, and testing sets containing 258, 65, and 57 samples, respectively.
 The training objective was to minimize the loss functions in the training set.
-We used AdamW optimizer with weight decay of $10^{-5}$ to optimize the loss functions.
+We used AdamW optimizer with a weight decay of $10^{-5}$ to optimize the loss functions.
 We adopted a cosine annealing scheduler to decay the learning rate from $2\times 10^{-4}$ to zero for 1000 epochs.
-The loss was first computed for each batch within an epoch and the overall loss for an epoch was calculated by taking a mean over all the batch losses.
+The loss was first computed for each batch within an epoch, and the overall loss for an epoch was calculated by taking a mean over all the batch losses.
 The model with the highest mean DSC in the validation phase was chosen for further evaluation of the test set.
-The code for the implementations is made available at: https://github.com/ObedDzik/pca_segment.git
+For SAM-Med3D, we evaluated both zero-shot performance and fine-tuning.
+Fine-tuning was conducted using the original loss function as well as comparative loss functions for evaluation.
+Training used a batch size of 2, an initial learning rate of $8\times 10^{-6}$ and 150 epochs.
+Optimization employed AdamW with a linear learning rate scheduler (decay factor 0.1).
 2.4 Loss Functions
-Dice Loss (DL):
-The Dice Loss maximizes the overlap between the predicted segmentation and the ground truth. In this work, for a specific cubic patch of an image, the Dice Loss for a batch was computed as,
+Dice Loss:
+The DL maximizes the overlap between the predicted segmentation and the ground truth. In this work, for a specific cubic patch of an image, the DL for a batch was computed as,
 $\mathcal{L}_{\text{Dice}}=1-\frac{1}{2}\sum_{c\in\{0,1\}}\frac{2\sum_{i}p_{i}(c)g_{i}(c)}{\sum_{i}p_{i}(c)+\sum_{i}g_{i}(c)+\epsilon}$ (1)
@@
 The Focal Loss [24] helps focus on hard-to-classify examples by down-weighting the easy ones. In this work, the Focal Loss for an image patch was computed over a batch as,
-$\mathcal{L}_{\text{Focal}}=-\frac{1}{2}\sum_{c\in\{0,1\}}\sum_{i}\alpha_{c}(1-p_{i}(c))^{\gamma}\log(p_{i}(c))$ (2)
-where $p_{i}$ is the predicted probability for a given class $c$ of a voxel $i$, $\alpha$ is a factor defining the balance between background and foreground classes in a binary segmentation task, and $\gamma$ is the focusing parameter that adjusts the rate at which easy examples are down-weighted.
-Dice Focal Loss (DFL):
-The Dice Focal Loss [25] combines the Dice Loss and the Focal Loss forming a compound loss function:
+$\mathcal{L}_{\text{Focal}}=-\sum_{c\in\{0,1\}}\sum_{i}g_{i}(c)\alpha_{c}(1-p_{i}(c))^{\gamma}\log(p_{i}(c))$ (2)
+where $\alpha$ is a factor defining the balance between background and foreground classes in a binary segmentation task, and $\gamma$ is the focusing parameter that adjusts the rate at which easy examples are down-weighted.
+Dice Focal Loss:
+The DFL [45] combines the DL and the Focal Loss forming a compound loss function:
 $\mathcal{L}_{\text{DFL}}=\mathcal{L}_{\text{Dice}}+\mathcal{L}_{\text{Focal}}$ (3)
-For the Dice Focal Loss formulation, we set $\gamma=2$ and $\alpha=1$ in all our experiments similar to the implementation reported in [26].
-L1-weighted Dice Focal Loss (L1DFL):
-The L1-weighted Dice Focal Loss (L1DFL) applies a weighting strategy to the Dice Loss (DL) based on the L1 norm between the predicted probabilities and the ground truth labels and combines it with the Focal Loss (FL).
-The Dice Loss generally has two variants, one with squared terms in the denominator [27] and one without [12].
-In L1DFL, we employ the Dice Loss with the squared denominator which was preferred over the non-squared version following its performance in [28].
+For the DFL formulation, we set the Focal Loss terms $\gamma=2$ and $\alpha=1$ in all our experiments similar to the implementation reported in [31].
+L1-weighted Dice Focal Loss:
+L1DFL applies a gradient harmonization weighting strategy to the Dice Loss based on both the L1 norm between predicted probabilities and ground truth labels (quantifying voxel-level difficulty) and the density of samples at each difficulty level (defined by bins), then combines this weighted Dice Loss with Focal Loss.
+The DL generally has two variants, one with squared terms in the denominator [29] and one without [8].
+In L1DFL, we employ the DL with the squared denominator which was preferred over the non-squared version following its performance in [26].
 We describe the mathematical formulation of the weighting strategy next.
@@
 where $n=\lceil{\frac{1}{\Gamma}}+1\rceil$ is the total number of bins.
-For instance, with a bin width of $\Gamma=0.1$ and a total of $n$ = 11 bins, $B_{k}$ takes values at regular intervals in the range [0,1], such that for $k=0,1,2,\dots,10$, bin centers ($B_{k}$) = $[0,0.1,0.2,\dots,1.0]$.
+For instance, with a bin width of $\Gamma=0.1$ ($n$ = 11 bins), $B_{k}$ takes values at regular intervals in the range [0,1], such that for $k=0,1,2,\dots,10$, bin centers ($B_{k}$) = $0,0.1,0.2,\dots,1.0]$.
 We then calculate the count of $\Delta$ values that fall within each bin:
@@
 $\mathcal{D}(B_{k})=\frac{C(B_{k})}{\lambda_{\kappa}(B_{k})}$ (9)
-The density $\mathcal{D}(B_{k})$ represents the concentration of samples around the bin center $B_{k}$.
-For $B_{k}$ values close to 0, the L1 norm values ($\Delta_{i}$) are also close to 00, indicating easy samples where the predictions align closely with the ground truth.
-In contrast, bins with centers near 1 represent L1 norms closer to 1, indicating hard-to-classify samples where predictions deviate significantly from the ground truth.
-Thus, based on the norm density, we can evaluate how common or rare a particular prediction difficulty is, allowing the weighting to be adjusted accordingly.
-This adjustment ensures that less frequent and difficult samples, which often correspond to foreground samples in an imbalanced scenario receive higher weights than the easy-to-classify samples.
-The weight for each bin is then calculated as:
+The density $\mathcal{D}(B_{k})$ quantifies how many voxels have L1 norms within the bin centered at $B_{k}$.
+The bin center represents the difficulty level: $B_{k}\approx 0$ indicates easy samples (predictions closely match ground truth), while $B_{k}\approx 1$ indicates hard samples (predictions deviate significantly from ground truth).
+Thus, density, $\mathcal{D}(B_{k})$, measures the prevalence of each difficulty level, and high density means many voxels share similar L1 norm values (commonly occurring difficulty), while low density means few voxels fall in that range.
+The weight for each bin is thus calculated as:
 $w(B_{k})=\frac{N}{\mathcal{D}(B_{k})}$ (10)
-These weights are applied to the numerator and the denominator of the Dice Loss to account for the imbalances caused by the variations in $\Delta$
+where $N$ is the total number of samples.
+Consequently, samples in high-density regions receive lower weights, while samples in low-density regions receive higher weights.
+This ensures that less frequent and difficult samples, which often correspond to small foreground structures in imbalanced scenarios, receive appropriate emphasis without allowing outliers or ambiguous examples to dominate optimization, unlike focal loss approaches that monotonically increase weights with difficulty.
+These harmonized weights are applied to both the numerator and denominator of the Dice loss to account for imbalances caused by variations in $\Delta$.
 $\mathcal{L}_{\text{wDice}}=1-\frac{2\sum_{i=1}^{N}w_{i}y_{i}p_{i}+\epsilon}{\sum_{i=1}^{N}w_{i}(y_{i}^{2}+p_{i}^{2})+\epsilon}$ (11)
@@
 If the L1 norms of the examples are uniformly distributed, the density $\mathcal{D}(B_{k})$ will be the same for all bins, resulting in equal weights $w(B_{k})$ for all bins.
-In this case, the weighted Dice Loss, ($L_{\text{wDice}}$), will reduce to the standard Dice Loss.
+In this case, the weighted DL, ($L_{\text{wDice}}$), will reduce to the standard DL.
 Finally, we combine $\mathcal{L}_{\text{wDice}}$ with the Focal Loss. The full expression of L1DFL thus is:
-[…]
-We illustrate the dynamic weighting strategy of L1DFL below.
-For a given 4x4 matrix, Figure 1 highlights how L1DFL downweights contributions from the background class and focuses on regions of classification difficulty.
-Specifically, the figure shows that higher weights are applied for voxels with L1 norm ($\Delta$) close to 1, but for voxels with L1 norms close to 0, representing that those voxels are easily classified, lower weights are applied.
-Thus, regions corresponding to false positives and negatives receive much attention from the model.
-In our experiments, we empirically selected a constant bin width, $\Gamma$, of 0.1 and $\gamma$ of 2 for the focal loss component of the loss function.
-Figure 1: An illustration of the dynamic weighting strategy in L1 weighted Dice Focal Loss (L1DFL).
-(a) shows the ground truth (G), where 1 indicates lesion areas and 0 indicates background.
-(b) displays the predicted probability map, where green regions denote false positive predictions and red regions represent ROIs.
-(c) shows the L1 norm values between the predicted probabilities and ground truth, highlighting the error magnitude for each voxel.
-Higher values imply higher classification difficulty.
-(d) presents the density, $\mathcal{D}$, which reflects the frequency distribution of each L1 norm value, capturing the difficulty of correctly classifying each voxel.
-Lower values imply higher classification difficulty.
-(e) depicts the final weights assigned to each voxel, where higher weights focus on rare, misclassified regions while down weighting well-classified background areas.
-This weighting strategy reduces the influence of abundant, correctly predicted voxels and increases the focus on difficult, misclassified lesions, thereby enhancing the model’s focus on challenging samples.
+$\text{L1DFL}=\mathcal{L}_{\text{wDice}}+\mathcal{L}_{\text{Focal}}$ (12)
+For our implementation, we empirically selected a constant bin width, $\Gamma$, of 0.1 and $\gamma=2$ for the Focal Loss component of L1DFL.
+We assess the robustness of the loss function to $\Gamma$ and $\gamma$ hyperparameters, computed over a five-epoch window centered at the epoch with the highest validation Dice.
+We report the performance in the results section (3.6).
+We illustrate the dynamic weighting strategy of L1DFL in Figure 1 using a 4×4 matrix.
+Figure 1: An illustration of the dynamic weighting strategy in L1DFL.(a) Ground truth (G) with lesion areas (1) and background (0).
+(b) Predicted probability map showing false positives (green) and regions of interest (red).
+(c) L1 norm values computed as $|\hat{y}-y|$ for each voxel, representing prediction error magnitude.
+Higher L1 norm values indicate greater classification difficulty.
+(d) Density $\mathcal{D}$, showing the frequency of each voxel’s difficulty level across all voxels.
+A difficulty level is defined by a range of L1 norm values (bin), and density measures how many voxels fall within that bin.
+(e) Final weights ($w=N/\mathcal{D}$) assigned through gradient harmonization.
+Voxels at high-density difficulty levels receive lower weights, while voxels at low-density difficulty levels receive higher weights.
+This prevents gradient concentration on any single difficulty level, distributing optimization effort more uniformly across the difficulty spectrum.
+2.5 Empirical Loss Analysis
+To empirically validate the gradient harmonization mechanism of L1DFL and compare it against baseline loss functions, we conducted two complementary analyses on converged models.
+Prediction Distribution Analysis:
+We evaluated the calibration properties and uncertainty quantification capabilities of each loss function by analyzing the distribution of prediction entropy and confidence scores.
+For each voxel in the validation set, we computed the prediction entropy as $H=-\sum_{c}p_{c}\log p_{c}$ where $p_{c}$ is the predicted probability for class $c$, and the prediction confidence as $\max_{c}p_{c}$.
+We stratified these distributions by prediction correctness (correct vs.
+incorrect predictions) to assess whether each loss function produces well-calibrated uncertainty estimates.
+The calibration gap, defined as the difference in mean confidence between correct and incorrect predictions, quantifies how well the model’s confidence aligns with its actual accuracy, a larger gap indicates better calibration.
+Sample Weighting Analysis:
+To directly demonstrate the gradient harmonization effect, we analyzed how each loss function weights samples of varying difficulty.
+For each test sample, we computed: (1) sample difficulty as the mean absolute difference between predictions and targets $d=|\hat{y}-y|$, representing the inherent difficulty of the prediction, and (2) effective sample weight as the $L_{2}$ norm of the loss gradient with respect to the model outputs $w=\|\nabla_{\hat{y}}\mathcal{L}\|_{2}$, representing how much each sample contributes to the gradient during optimization.
+We computed the Pearson correlation coefficient $\rho(d,w)$ between difficulty and weight across all samples.
+A lower correlation indicates that the loss function assigns more balanced weights across the difficulty spectrum, preventing gradient concentration on outliers.
+We also visualized the relationship between difficulty and weight by binning samples into 50 difficulty levels and computing mean weights per sample for each bin.
+This analysis reveals whether a loss function over-emphasizes hard samples (high $\rho$) or achieves balanced weighting preventing outlier dominance.
+Both analyses were performed on a converged SegResNet model trained with the different loss functions using the validation set, with sliding window inference for computational efficiency.
 2.6 Model Evaluation
+We evaluated the performance of the loss functions on the test set using patient- and lesion-level metrics.
+Predictions were generated using a sliding-window inference strategy and resampled to the original image coordinates.
+In addition to overall test-set performance, results were stratified by lesion burden into single-lesion and multiple-lesion scenarios, defined as images containing exactly one lesion and two or more lesions, respectively.
+For each scenario and for the overall test set, we performed both patient-level and lesion-level evaluations.
+At the patient level, we assessed segmentation accuracy using the DSC, detection performance using true positives (TP), false positives (FP), false negatives (FN), and the F1 score, and clinically motivated metrics based on TMTV.
+For multiple-lesion cases, we additionally evaluated performance with respect to lesion spatial dissemination ($D_{\text{max}}$).
+Lesion-level evaluation focused on segmentation accuracy (DSC) and performance across different molecular tumor volume (MTV) ranges of individual lesions.
+Detailed mathematical definitions, matching strategies, thresholding procedures, and clinical metric formulations are provided in B.
+3 Results
+3.1 Segmentation performance across different networks
+In this section, we present results for the segmentation performance of the three loss functions, DL, DFL, and L1DFL across the different architectures (Table 1).
+We report the average and median DSCs at the patient level, with mean values of true positive, false positive, and false negative counts on the test set and the F1 score performance.
+A one-tailed paired Wilcoxon signed-rank test with Bonferroni correction was performed to evaluate the significance of the performance of L1DFL from those of DL and DFL separately, with corresponding p-values reported.
+Network Loss Function DSC Mean Rates F1 scores P values
+Mean Median TP FP FN
+Attention U-Net DL 0.54 $\pm$ 0.24 0.58 [0.44, 0.72] 0.81 $\pm$ 0.35 2.06 $\pm$ 1.89 0.19 $\pm$ 0.35 0.50 $\pm$ 0.28 $0.004$
+DFL 0.51 $\pm$ 0.26 0.54 [0.33, 0.73] 0.77 $\pm$ 0.36 2.65 $\pm$ 2.60 0.23 $\pm$ 0.36 0.44 $\pm$ 0.27 <0.001<0.001*
+L1DFL 0.58 $\pm$ 0.27 0.66 [0.51, 0.77] 0.77 $\pm$ 0.37 0.42 $\pm$ 0.65 0.23 $\pm$ 0.37 0.69 $\pm$ 0.34
+SegResNet DL 0.52 $\pm$ 0.30 0.60 [0.29, 0.76] 0.73 $\pm$ 0.39 0.73 $\pm$ 1.10 0.27 $\pm$ 0.39 0.62 $\pm$ 0.35 0.235
+DFL 0.52 $\pm$ 0.25 0.59 [0.41, 0.71] 0.81 $\pm$ 0.34 2.05 $\pm$ 2.00 0.19 $\pm$ 0.34 0.49 $\pm$ 0.27 0.015
+L1DFL 0.57 $\pm$ 0.29 0.68 [0.47, 0.78] 0.76 $\pm$ 0.37 0.52 $\pm$ 0.76 0.24 $\pm$ 0.37 0.66 $\pm$ 0.34
+U-Net DL 0.49 $\pm$ 0.28 0.55 [0.28, 0.75] 0.69 $\pm$ 0.38 1.27 $\pm$ 1.18 0.30 $\pm$ 0.39 0.49 $\pm$ 0.29 0.197
+DFL 0.49 $\pm$ 0.28 0.54 [0.25, 0.70] 0.71 $\pm$ 0.39 1.66 $\pm$ 1.53 0.28 $\pm$ 0.40 0.46 $\pm$ 0.29 0.165
+L1DFL 0.52 $\pm$ 0.29 0.60 [0.38, 0.76] 0.71 $\pm$ 0.38 0.71 $\pm$ 0.79 0.29 $\pm$ 0.39 0.58 $\pm$ 0.33
+UNETR DL 0.37 $\pm$ 0.26 0.39 [0.11, 0.59] 0.64 $\pm$ 0.43 7.74 $\pm$ 7.70 0.33 $\pm$ 0.42 0.21 $\pm$ 0.19 0.435
+DFL 0.39 $\pm$ 0.28 0.46 [0.14, 0.61] 0.64 $\pm$ 0.42 5.49 $\pm$ 5.77 0.32 $\pm$ 0.40 0.26 $\pm$ 0.21 0.518
+L1DFL 0.39 $\pm$ 0.29 0.48 [0.08, 0.62] 0.61 $\pm$ 0.41 2.59 $\pm$ 4.11 0.35 $\pm$ 0.42 0.38 $\pm$ 0.29
+SAM-Med3D Zero-shot 0.31 $\pm$ 0.24 0.26 [0.11, 0.45] 0.78 $\pm$ 0.35 2.32 $\pm$ 3.39 0.33 $\pm$ 0.46 0.53 $\pm$ 0.36 <0.001<0.001*
+Finetuned (DCE) 0.51 $\pm$ 0.25 0.51 [0.29, 0.67] 0.74 $\pm$ 0.37 0.45 $\pm$ 1.05 0.29 $\pm$ 0.38 0.71 $\pm$ 0.37 0.712
+DL 0.48 $\pm$ 0.26 0.51 [0.32, 0.67] 0.75 $\pm$ 0.37 0.90 $\pm$ 3.40 0.31 $\pm$ 0.45 0.70 $\pm$ 0.38 0.290
+DFL 0.48 $\pm$ 0.26 0.51 [0.30, 0.68] 0.73 $\pm$ 0.37 0.64 $\pm$ 2.24 0.29 $\pm$ 0.39 0.71 $\pm$ 0.38 0.254
+L1DFL 0.49 $\pm$ 0.26 0.50 [0.25, 0.74] 0.78 $\pm$ 0.34 1.55 $\pm$ 3.56 0.33 $\pm$ 0.49 0.69 $\pm$ 0.39
+Table 1: Comparison of the loss functions on the test set based on mean and median patient-level DSC.
+The mean DSC, TP, FP, FN and F1 scores are presented with standard deviations and the median DSC with interquartile ranges.
+A detection is considered a true positive if the predicted mask overlaps with the voxel containing the SUVmax value in the ground truth.
+The P values reflect a one-tailed paired Wilcoxon signed-rank test comparing the median difference in DSC between L1DFL and DL and DFL, with * indicating statistical significance at the $\alpha=0.003$ level (Bonferroni correction applied).
+The CNN-based models generally outperformed the transformer-based network and the foundation model, likely due to the limited training data and the inherent inductive bias of CNNs.
+Across all architectures, median DSC values were consistently higher than the mean DSCs.
+UNETR exhibited the poorest overall performance, with the DL yielding the lowest DSC and F1 scores, despite achieving competitive mean TP counts.
+The low DSC and F1 were primarily due to a high FP rate.
+Similarly, SAM-Med3D showed the lowest DSC in zero-shot testing, although TP rates remained competitive; again, poor performance was linked to elevated FP rates.
+When fine-tuning SAM-Med3D with different loss functions, including the original Dice Cross-Entropy loss (DCE), median DSCs were similar, with L1DFL lagging by 1 point.
+However, L1DFL achieved the highest TP counts.
+Notably, although SAM-Med3D is unimodal and received only PET input, its performance was competitive with multimodal CNN models, likely benefiting from features learned during pretraining.
+Within the CNN models, L1DFL consistently achieved the best balance of DSC, FP rate, and F1 score.
+The combination of L1DFL with SegResNet yielded the highest median DSC $(0.68)$, whereas the best mean TP counts were obtained by DL and DFL on SegResNet and Attention U-Net, respectively.
+However, these high sensitivity values came at the cost of increased false positives, reducing F1 scores.
+For example, DL on SegResNet had an FP rate of $0.73\pm 1.10$, whereas on Attention U-Net it had $2.06\pm 1.89$, roughly three times higher.
+Overall, Attention U-Net and SegResNet outperformed U-Net.
+DL consistently achieved higher median DSC, FP, and F1 scores than DFL across all CNNs, although DFL had higher TP counts on SegResNet and U-Net.
+For the transformer-based model, DFL slightly outperformed DL in FP and F1 scores.
+L1DFL, however, achieved the best balance of TP detection and minimal FPs on both CNNs and UNETR, and showed higher median DSC in the third quartile range for all models.
+For instance, on UNETR where DL and DFL had mean FP rates above 5.0, L1DFL minimized false detections while maintaining lesion sensitivity, yielding the highest mean and median DSCs for this network.
+Although a one-tailed Wilcoxon signed-rank test comparing loss functions across networks did not provide sufficient evidence to reject the null hypothesis, statistically significant differences were observed on Attention U-Net when comparing L1DFL to DFL.
+The training and validation dynamics of all networks and loss functions are provided in A, showing consistent convergence across models.
+Figure 2: Segmentation results on a test PET image.
+The first column shows the coronal maximum intensity projection (MIP), ground truth (GT) mask, and GT overlaid on the PET image.
+Columns correspond to different networks, and rows correspond to different loss functions.
+Figure 2 illustrates the segmentation performance of the different loss functions on a test PET image.
+In this example, SegResNet achieved the most accurate lesion segmentation with the highest DSC among all models.
+DL generally outperformed DFL, except for SegResNet.
+Across all networks, L1DFL provided the best performance, including on the transformer-based models where DL and DFL struggled.
+3.2 Performance in single and multiple lesion scenarios
+We evaluated the loss functions in two lesion scenarios: single lesion and multiple lesions in an image.
+Figure 3 shows that DSC values were more widely spread in the single-lesion scenario across all networks.
+The largest interquartile range (IQR) differences between scenarios were observed for UNETR across all loss functions and SegResNet with DL, whereas Attention U-Net and SAM-Med3D maintained tighter IQRs.
+Additionally, DSC values beyond the 75th percentile were generally higher in the single-lesion scenario than the multiple lesion scenario, with more cases exceeding $0.75$.
+For the single lesion scenario, L1DFL performed consistently better than the other loss functions.
+Specifically, on the CNN models, L1DFL improved median DSC by at least $19\%$ over DL and DFL, while for UNETR and SAM-Med3D, improvements were at least $45\%$ and $3.3\%$, respectively.
+In contrast, DL and DFL showed variable performance: DL outperformed DFL on Attention U-Net, SegResNet, and SAM-Med3D, while DFL achieved higher median DSC on U-Net and UNETR.
+L1DFL’s best performance was on SegResNet with a median DSC of $0.69$, whereas DL and DFL performed best on SAM-Med3D with a median DSC of $0.64$ and $0.62$, respectively.
+In the multiple-lesion scenario, even though DL generally achieved the best overall performance across the networks, L1DFL had the highest median DSCs on UNETR and Attention U-Net.
+Conversely, SAM-Med3D struggled to achieve accurate lesion segmentation in this scenario, and although its overall median DSC (Table 1) was higher than UNETR, UNETR performed better at segmenting multiple lesions.
+For SAM-Med3D, all loss functions had median DSCs in the range $[0.22,0.25]$, with tighter IQRs suggesting generally poor performance on cases with multiple lesions in the image.
+Figure 3: Box plots overlaid with swarm plots illustrating the distribution of Dice Similarity Coefficient (DSC) performance for the three loss functions across two lesion scenarios: Multiple lesions (lesion count >2>2) and Single lesion (lesion count =1=1).
+Each box plot’s top and bottom edges represent the interquartile range (IQR), while the horizontal lines within the boxes indicate the median DSC values.
+Whisker lengths are set to 1.5 times the IQR.
+The number of cases in the single lesion scenario was N = 34 and for the multiple lesion scenario, the number of cases was N = 23.
+3.3 Performance based on lesion molecular volume
+We also assessed how the loss functions performed across varying tumor volumes in both single- and multiple-lesion scenarios (Figure 4).
+DL yielded higher DSCs for smaller lesions, whereas DFL was slightly more robust to lesion size variations, especially on CNN-based models.
+Nevertheless, performance for both loss functions generally declined with increasing tumor volume.
+For instance, in single-lesion scenarios, DL and DFL achieved higher DSCs at lower volumes, but their performance decreased more sharply at larger volumes, occasionally resulting in relatively better outcomes for multiple-lesion cases.
+Besides, the segmentation accuracy for both loss functions was higher at the lesion-wise assessment level than at the patient level for smaller tumors, likely due to the greater impact of false positives.
+L1DFL, in contrast, demonstrated a consistent advantage over DL and DFL, particularly at the patient level, where it exhibited an upward DSC trend with increasing TMTV across all networks and remained stable across different lesion scenarios.
+Moreover, on the CNN-based models, L1DFL consistently maintained DSC values between $0.6$ and $0.8$ across both patient- and lesion-level assessments, whereas DL and DFL either started within this range and declined as volume increased or began lower and gradually rose into it.
+From a network perspective, CNNs generally achieved higher segmentation performance than UNETR and SAM-Med3D, with Attention U-Net and SegResNet performing best among the CNNs, and U-Net performing less favorably.
+DL paired well with SegResNet, while DFL produced similar results on SegResNet and Attention U-Net.
+Both losses, however, showed reduced performance on U-Net, particularly for larger lesions; for example, in the single-lesion scenario beyond an MTV threshold of $5$ mL, DL’s median DSC decreased from $0.73$ to $0.32$, and DFL from $0.81$ to $0.48$.
+L1DFL, on the other hand, remained relatively stable, with only a $14\%$ decrease.
+On Attention U-Net, DL decreased by $66\%$ at high MTV values and DFL by $21\%$, while L1DFL exhibited only a slight dip with a mild upward trend overall.
+UNETR and SAM-Med3D exhibited more variable performance across assessment levels.
+On UNETR, DL and DFL declined in performance with increasing tumor volumes, whereas L1DFL remained robust, showing a downward trend only in lesion-wise analysis for single-lesion cases.
+SAM-Med3D performed better for single-lesion cases than UNETR, although DL and DFL still exhibited decreasing performance.
+In contrast, for multiple-lesion cases, SAM-Med3D showed a sharp drop in segmentation accuracy.
+Overall, finetuning SAM-Med3D with all three loss functions did not improve performance on the multiple lesions cases, with L1DFL trailing DL and DFL by at least $40\%$.
+Figure 4: Patient- and lesion-level median Dice Similarity Coefficient (DSC) performance of the loss functions across varying ground truth tumor volumes.
+Each column corresponds to a network, with subcolumns showing single-lesion (S) and multiple-lesion (M) scenarios.
+The top row presents patient-level results versus total molecular tumor volume (TMTV), while the bottom row shows lesion-level results versus molecular tumor volume (MTV).
+Thresholds $(t)$ span the interquartile range up to the 85th percentile, and median DSC is computed for all lesions $(l)$ with volume $(v_{l}>t)$.
+Line colors indicate different loss functions.
+3.4 Performance based on lesion dissemination
+To quantify the spatial distribution of lesions, we calculated $D_{\text{max}}$, the maximum Euclidean distance between any two lesions in the image.
+For the performance assessment of the loss functions based on the $D_{\text{max}}$, we categorized the distances into the following groups: 0-9 cm (G0), 9-11 cm (G1), 11-14 cm (G2), and 14-60 cm (G3), representing the first, second, third and fourth quartile ranges, respectively.
+Patient-level analysis of DSC showed consistent improvement across these ranges (Figure 5) with differences in segmentation performance across architectures.
+CNN-based models generally performed better on widely spread lesions (G2 and G3) than on closer distances, although performance declined from G2 to G3, indicating increasing challenges as tumors extend further.
+While SAM-Med3D and UNETR performed worse overall than the CNN-based models, they showed relatively stronger performance at lower $D_{\text{max}}$ ranges, with accuracy decreasing as lesions spread further, even in the presence of higher tumor activity indicated by the TLA plots.
+Moreover, performance generally varied with the choice of loss function for each network.
+At lower distances (G0 and G1), DL with SegResNet achieved the highest median DSCs, while L1DFL paired with Attention U-Net offered greater consistency.
+At higher distances (G2 and G3), L1DFL with Attention U-Net or SegResNet provided the strongest and most consistent results among the CNN models, with the widest IQR being $0.17$.
+Additionally, U-Net with L1DFL was competitive at G2, reaching a median DSC of $0.74$ and outperforming DL and DFL on all three CNN models by at least $5\%$.
+DL, on the other hand, performed best in the first and fourth quartiles on U-Net, while DFL yielded improved performance in the second and third quartiles.
+Overall, Attention U-Net with L1DFL provided the most consistent performance across $D_{\text{max}}$ ranges while SegResNet was most robust at larger lesion spreads.
+U-Net performed best at lower distances.
+On SAM-Med3D, all loss functions showed reduced performance, with median DSCs ranging from $0.20\text{ – }0.40$ at lower lesion spreads (G0 and G1) and dropping to $0.10\text{ – }0.25$ at higher spreads (G2 and G3).
+Despite the generally poor results, DL and DFL achieved a $40\%$ improvement over UNETR in the second quartile, although their performance remained below CNN-based models.
+In contrast, L1DFL performed better when paired with UNETR than with SAM-Med3D, albeit with less consistency as lesion spread increased, as indicated by the wider IQR.
+DFL showed the highest performance on UNETR in the first and last quartiles, while L1DFL performed best at G1 and G2.
+Figure 5: Plots illustrating the Dice Similarity Coefficients (DSC) across four interquartile ranges of $D_{\text{max}}$.
+The groups G0 (0-9 cm), G1 (9-11cm), G2 (11-14cm), and G3 (14-60cm) correspond to the interquartile ranges of $0-25\%$, $25-50\%$, $50-75\%$, and $75-100\%$, respectively.
+Each column show performance on a specific network.
+In the box plots, horizontal lines indicate the median, and whisker lengths are set to 1.5 times the interquartile range (IQR), with outliers represented as black dots.
+Accompanying bar plots in the last column display the median total lesion activity (TLA) values for cases within each $D_{\text{max}}$ group.
+The different loss functions are color-coded for clarity.
+The total number of cases in each group was N = 6.
+3.5 Loss Function Properties
+3.5.1 Prediction Distribution Analysis
+From Table 2, L1DFL demonstrates better calibration than the baseline methods.
+Its calibration gap (0.2859) is substantially larger than that of DFL (0.0263) and DL (0.0129), indicating stronger separation between correct and incorrect predictions.
+When errors occur, L1DFL appropriately expresses uncertainty, with a mean entropy of 0.4762 for incorrect predictions versus 0.0212 for correct ones.
+In contrast, DFL and DL remain highly overconfident even on incorrect predictions, with confidence scores of 0.9737 and 0.9871, respectively.
+Table 2: Quantitative comparison of loss function properties
+Prediction Distribution Sample Weighting
+Loss Function Entropy (correct) Entropy (incorrect) Calibration Gap Confidence (correct) Confidence (incorrect) Difficulty-Weight Correlation
+L1DFL 0.0212 0.4762 0.2859 0.9982 0.7123 0.3376
+DFL 0.0000 0.0614 0.0263 1.0000 0.9737 0.9587
+DL 0.0000 0.0304 0.0129 1.0000 0.9871 0.0961
+These trends are reflected in the entropy distributions (Figure 6, top row).
+L1DFL exhibits a clear bimodal structure, with correct predictions concentrated at near-zero entropy and incorrect predictions at higher entropy values, indicating the model appropriately recognizes ambiguous cases.
+Conversely, DFL and DL collapse to near-zero entropy for both correct and incorrect predictions, demonstrating poor uncertainty discrimination.
+Similarly, the confidence distributions (Figure 6, bottom row) show that L1DFL assigns lower confidence to incorrect predictions while maintaining high confidence for correct ones.
+In contrast, DFL and DL produce near-saturated confidence values for all predictions, offering little discriminative value for uncertainty estimation.
+Figure 6: Entropy (top) and confidence (bottom) distributions for correct and incorrect predictions.
+The model trained with L1DFL exhibits clear separation between confident correct predictions and uncertain incorrect predictions, whereas DFL and DL remain overconfident regardless of prediction correctness.
+Figure 7 illustrates the L1-norm distribution of a converged SegResNet trained with L1DFL.
+As expected, the fraction of easy examples is large and could dominate the overall gradient.
+However, a non-negligible number of very difficult examples remain misclassified.
+As noted by Li et al.
+[22], such samples can be treated as outliers, and forcing the model to fit them may degrade performance on the majority of data due to their inconsistent gradient directions.
+In addition, a substantial fraction of samples lies near the decision boundary, reflecting intrinsic ambiguity or class overlap.
+The prominence of this ambiguous region indicates that the model captures uncertainty rather than enforcing overconfident predictions, consistent with the bimodal entropy and confidence distributions shown in Figure 6.
+Figure 7: L1-norm distribution of a converged model trained with L1DFL (log-scaled y-axis).
+3.5.2 Sample Weighting Analysis
+Figure 8 provides direct evidence of L1DFL’s gradient harmonization mechanism through analysis of effective sample weighting across difficulty levels.
+The difficulty-weight correlation for L1DFL $(0.3376)$ is substantially lower than DFL $(0.9587)$ indicating that L1DFL assigns more uniform weights across the difficulty spectrum rather than exhibiting strong monotonic dependence on sample difficulty.
+DL on the other hand has the lowest correlation of $0.0961$ indicating a difficulty agnostic behavior in weighting.
+Figure 8 shows that L1DFL assigns relatively balanced weights across difficulty levels, with a controlled emphasis on moderately difficult samples in the 0.60.60.80.8 range, consistent with its intended gradient harmonization behavior.
+In contrast, DFL applies monotonically increasing weights, with gradient magnitudes growing rapidly as difficulty approaches 1.0, causing optimization to be dominated by the hardest samples, including outliers and ambiguous cases.
+DL exhibits more irregular weighting patterns, lacking a clear and consistent structure across difficulty levels.
+Figure 8: Comparison of sample weighting across difficulty levels for the different loss functions.
+L1DFL emphasizes moderately difficult samples, whereas DFL heavily weights the hardest examples and DL shows inconsistent weighting.
+3.6 Ablation Study and Hyperparameter Robustness Test
+To assess the contribution of each component of the proposed L1DFL loss, we conducted an ablation study on two representative architectures, SegResNet and UNETR.
+Starting from a Dice+Focal baseline, we progressively modified the voxel-wise weighting strategy by (i) replacing the histogram-based weighting with raw gradient magnitudes, (ii) using squared gradients, and (iii) applying the full proposed formulation.
+All experiments were performed using the same training protocol and hyperparameters.
+Table 3: Ablation study evaluating the impact of individual components of the proposed L1DFL loss on validation Dice score (mean $\pm$ std).
+Loss Variant SegResNet UNETR
+Dice + Focal (baseline) 0.50 $\pm$ 0.02 0.33 $\pm$ 0.02
+Only gradients ($w=\Delta$) 0.50 $\pm$ 0.02 0.35 $\pm$ 0.02
+Gradients squared ($w=\Delta^{2}$) 0.47 $\pm$ 0.03 0.33 $\pm$ 0.02
+L1DFL (proposed) 0.53 $\pm$ 0.02 0.36 $\pm$ 0.02
+As shown in Table LABEL:tab:ablation, using raw gradients provides a small improvement over the baseline for UNETR but has little effect for SegResNet.
+Squaring the gradients ($w=\Delta^{2}$) reduces performance for both architectures, highlighting the importance of controlled weighting.
+The full L1DFL formulation achieves the highest Dice scores for both networks, demonstrating that each component, including the histogram-based gradient weighting and the L1 normalization, contributes to improved segmentation performance.
+We conducted a validation-based hyperparameter sensitivity analysis.
+Table 4 reports the mean and standard deviation of the validation Dice across a range of $\Gamma$ and $\gamma$ values.
+Performance varies slightly across configurations, ranging from $0.51$ to $0.54$, with overlapping standard deviations, indicating that L1DFL is robust to these hyperparameter settings.
+We present the results using the SegResNet model; however, comparable ranges of Dice score differences were also observed with SAM-Med3D and UNETR.
+Table 4: Validation-set Dice scores for different $\gamma$ (Focal Loss) and bin width $\Gamma$ values.
+Values are reported as mean $\pm$ standard deviation over a 5-epoch window centered at the best validation epoch.
+$\gamma$ $\Gamma$ Dice
+1 0.05 0.53 $\pm$ 0.03
+1 0.10 0.51 $\pm$ 0.02
+1 0.20 0.51 $\pm$ 0.04
+2 0.05 0.53 $\pm$ 0.04
+2 0.10 0.53 $\pm$ 0.02
+2 0.20 0.54 $\pm$ 0.03
+5 0.05 0.53 $\pm$ 0.03
+5 0.10 0.53 $\pm$ 0.02
+5 0.20 0.54 $\pm$ 0.01
+4 Discussion
+In this work, we proposed and evaluated a novel loss function L1DFL, in segmenting metastatic prostate cancer lesions based on different lesion scenarios characterized by the number of lesions in an image.
+Specifically, we defined two scenarios: single lesion scenario which include whole body PET/CT images with only a single lesion present in the ground truth mask and multiple lesion scenario being images with more than one lesion.
+We compared the performance of our proposed loss function with the DL and DFL.
+Moreover, we report how the size of a lesion and its spatial distribution influence the segmentation result.
+The dataset used in this study included tumors which had resurfaced after treatment with relatively low molecular volumes and standardized uptake values, thus introducing unique challenges for correct segmentations.
+The CNN-based models achieved better overall segmentation performance, with higher true positive detections and median DSCs compared to UNETR and SAM-Med3D.
+This outcome is expected, as the limited training data favored CNNs, whose inductive bias enables them to learn relevant features more effectively from smaller datasets than transformer-based models [27, 20].
+Nevertheless, SAM-Med3D, as a pretrained medical foundation model, demonstrated competitive results when fine-tuned on the limited dataset.
+Our findings further showed that the performance of the loss functions is not uniform across different architectures.
+For instance, the results show that the DL generated more false positives on Attention U-Net compared to SegResNet, despite showing higher sensitivity on the former.
+Similarly, the DFL exhibited lower precision on the Attention U-Net architecture due to increased false positives.
+Both loss functions also had lower F1 scores on U-Net suggesting a lower balance of true positives and false positives.
+By evaluating performance at both the patient and lesion-levels, the effect of false positive rates was highlighted.
+There was a more than $10\%$ decrease in the median DSC values at the lesion level for both loss functions when assessments were made at the patient level, especially for lower tumor volume thresholds.
+L1DFL, in contrast, was more robust to these false detections.
+Additionally, by assessing performance based on the different lesion scenarios, the influence of various lesion characteristics such as volume, extent of spread, number of lesions, on accuracy could be evaluated.
+For instance, considering the single lesion scenario, the performance of the DL is shown to decline as the threshold volume of tumors increased beyond $5\text{ml}$.
+This could suggest that DL is highly sensitive to the voxels of high uptake values [40].
+This observation aligns with claims in [25] where it is indicated that the DL is inherently biased toward smaller regions and may generally perform poorly on larger lesions with more variable SUV distribution across the volume.
+The DL is based on overlap measurement and it inherently balances precision with recall, making it more robust in scenarios with inter-class class imbalances as small objects tend to acquire greater contribution to the overall loss [35, 25].
+However, it does not distinguish between intra-class imbalances leading to a balanced focus on both easy and difficult examples [18, 43].
+These two imbalance scenarios tend to dominate among recurrent prostate cancer images, consequently, at lower tumor volumes, the DL had higher DSC values than the DFL, yet struggles to maintain a consistently higher performance on scenarios of multiple lesions with larger volumes.
+This trend, however, is not limited to the DL, as slight performance declines were observed in the multiple-lesion setting, particularly at the patient assessment level for all losses, indicating added complexity as lesion count increases.
+In fact, with more lesions, different anatomical sites may be involved, and the effect of other organs with high uptake values could be pronounced [40].
+Besides, the evaluation based on lesion spread, $D_{\text{max}}$, highlights the decrease in the models’ performance as distribution across the body becomes more extensive.
+Overall, L1DFL produced the most robust segmentations across evaluation scenarios.
+Its superior performance stems from a gradient harmonization mechanism that mitigates two central challenges in imbalanced medical image segmentation: dominance of easy background voxels and instability caused by outlier samples.
+By estimating mini-batch statistics of sample difficulty (via the L1 norm) and difficulty density, L1DFL redistributes gradient contributions such that frequent difficulty levels are down-weighted while underrepresented ones are emphasized.
+This prevents training from being dominated by either abundant background voxels or extreme cases [22].
+Our analysis showed that extreme-difficulty samples (L1 norms of 0.90.91.01.0) persist after convergence and likely correspond to outliers, ambiguous annotations, or imaging artifacts.
+These samples occur more frequently than moderately difficult ones and may introduce inconsistent gradient directions, destabilizing optimization [22].
+By down-weighting them according to their prevalence, L1DFL reduces sensitivity to noisy labels and spurious correlations.
+Empirically, L1DFL also yields a bimodal prediction-entropy distribution, clearly separating confident correct predictions from appropriately uncertain ones.
+This behavior is particularly valuable in clinical settings where reliable uncertainty estimation is essential [2].
+In contrast, although the DL weights samples equally irrespective of the difficulty level [43] and the Focal Loss encourages the model to focus more on hard samples [24], the DFL tends to lose the DL component [41], thus, assigns monotonically increasing weights with difficulty, overemphasizing extreme cases regardless of prevalence, leading to overconfident errors.
+Moreover, the proposed loss function introduces additional hyperparameters ($\gamma$ for focal loss, $\Gamma$ for bin width) that require tuning.
+While we show robustness to these parameters in this study, optimal values may vary across datasets and applications.
+Also, L1DFL incurs higher computational cost than standard losses due to computing L1 norms, binning samples, calculating density, and applying density-based weights.
+While this overhead remains practical for medical image segmentation tasks in our experiments, it may be non-negligible for extremely large 3D volumes.
+The benefits of gradient harmonization are most pronounced in highly imbalanced scenarios with diverse difficulty distributions.
+For a relatively balanced distribution, the L1DFL reduces to the Dice Focal loss with no added advantage to the additional complexity.
+Additional limitations of this study include the dataset being restricted to patients with up to five PSMA-avid lesions, representing a consecutive oligometastatic BCR population; thus, the performance of the loss functions in more extensive disease remains untested.
+Although improvements were observed across evaluation metrics, most did not reach statistical significance.
+Inter-observer variability in the ground truth annotations was not assessed, so segmentation errors may have influenced model performance.
+Finally, the generalizability of L1DFL to other medical imaging modalities remains uncertain, as the analysis was limited to PET/CT scans.
+5 Conclusion
+We evaluated the performance of a novel loss function, L1DFL, in the detection and segmentation of metastatic prostate cancer lesions in PSMA PET/CT scans across single and multiple lesion scenarios, characterized by lesion count.
+L1DFL employs a gradient harmonization scheme that leverages sample difficulty, quantified by L1 norms, and the density of difficulty levels to modulate the Dice Loss, which is then combined with Focal Loss to emphasize challenging voxel classifications.
+We compared L1DFL against DL and DFL across three 3D CNN architectures (Attention U-Net, SegResNet, and U-Net), a transformer-based model (UNETR), and a medical foundation model (SAM-Med3D).
+While DL generally showed higher sensitivity to small lesions, both DL and DFL exhibited variable performance on larger volumes and more extensive lesion spread, with DFL in particular producing higher false positive rates.
+In contrast, L1DFL led to balanced segmentation performance across lesion scenarios and architectures, achieving higher F1 scores and Dice Similarity Coefficients.
+Future work will be directed at exploring the generalizability of L1DFL to other medical imaging datasets and tasks.
+6 Acknowledgments
+This work was supported by the Canadian Institutes of Health Research (CIHR) Project Grants PJT-162216 and PJT-173231.
+We also acknowledge computational resources and services provided by Microsoft AI for Health.
+7 Ethics Statement
+The datasets used in this work included 18F-DCFPyL PET/CT scans of patients diagnosed with biochemical recurrence of prostate cancer as part of an ongoing clinical trial (ClinicalTrials.gov NCT02899312) with informed written consent obtained from the participants.
+The ethics approval was granted by the UBC BC Cancer Research Ethics Board (REB) (REB Number: H16-01551).
+8 Declaration of Competing Interest
+The authors declare that they have no known competing personal or financial interests that could have influenced this work.
+9 Declaration of generative AI and AI-assisted technologies in the writing process
+During the preparation of this work the author used OpenAI GPT o1 in order to refine statements.
+After using this tool, the author reviewed and edited the content as needed and takes full responsibility for the content of the publication.
+10 Data and Code Availability
+The dataset used in this study is proprietary and cannot be publicly released due to data privacy restrictions.
+However, the implementation code used to develop and evaluate the proposed method is publicly available at: https://github.com/ObedDzik/pca_segment.git.
+Appendix A Training Behavior Across Networks and Losses
+Figure 9 presents the training loss (top row) and validation DSC (bottom row) curves across networks and loss functions.
+Across all models, the training losses decreased consistently, confirming stable convergence.
+Attention U-Net, SegResNet, and U-Net exhibited rapid early loss reduction, with DL and DFL converging toward the lowest final loss values.
+In contrast, L1DFL maintained a higher residual loss but plateaued steadily, suggesting a different optimization dynamic.
+UNETR followed a similar pattern, albeit with slower convergence.
+SAM-Med3D showed higher apparent loss values.
+Validation DSC curves highlighted differences in generalization.
+L1DFL consistently yielded the highest DSC values across all networks, reaching approximately 0.550.550.580.58, and showing stable trajectories with minimal late-epoch decline.
+DFL provided intermediate performance, while DL typically converged more slowly and plateaued at lower DSC values.
+Interestingly, U-Net and Attention U-Net showed sharp early gains under L1DFL, while UNETR’s DSC curves lagged behind, indicating slower but eventually comparable improvements.
+For SAM-Med3D, DSC curves converged more quickly and stabilized by 20 epochs, reflecting its pretraining advantage.
+Figure 9: Training and validation curves across five segmentation networks and four loss functions, DL, DFL, L1DFL, and DCE (the original loss function for SAM-Med3D).
+The top row shows training loss trajectories, while the bottom row shows validation Dice Similarity Coefficient evolution over epochs.
+Columns correspond to different network architectures and loss functions are color coded.
+SAM-Med3D training loss values were rescaled by a factor of 10 for visualization.
+Appendix B Model Evaluation
 We assessed the performance of the loss functions on the overall test data.
-We made predictions on the PET/CT whole-body images using the sliding-window technique [29] with a window of dimension (128, 128, 128) for all networks.
+We made predictions on the PET/CT (only PET for SAM-Med3D) whole-body images using the sliding-window technique [4] with a window of dimension (128, 128, 128) for all networks.
 The test set predictions were resampled to the coordinates of the original ground truth masks to compute the evaluation metrics.
@@
 Then, for each scenario, as well as on the overall test set, we performed patient-level and lesion-level assessments based on the methodologies below.
-2.7.1 Patient-Level Analysis
+B.0.1 Patient-Level Analysis
 Segmentation Metrics:
@@
 The F1 score is thus computed as:
-[…]
+$\text{F1}=\frac{\text{TP}}{\text{TP}+\frac{1}{2}(\text{FP}+\text{FN})}$ (14)
 Figure 10: Illustration for defining a true positive detection based on an overlap with the voxel containing the maximum standardized uptake value (SUVmax) in the ground truth lesion.
@@
 We categorized the calculated distances into inter-quartile ranges (IQR) indicating the first, second, third and fourth IQR and analyzed the performance of the loss functions on each group.
-2.7.2 Lesion-Level Analysis
+B.0.2 Lesion-Level Analysis
 Segmentation Metrics:
@@
 $\text{DSC}_{l}=\frac{2|G_{l}\cap P_{l}|}{|G_{l}|+|P_{l}|}$ (16)
-where $G_{l}$ and $P_{l}$ denote the ground truth and predicted masks of a specific lesion. For a given ground truth lesion with no match, $|G_{l}\cap P_{l}|=0$ yielding a DSC of 00.
+where $G_{l}$ and $P_{l}$ denote the ground truth and predicted masks of a specific lesion. For a given ground truth lesion with no match, $|G_{l}\cap P_{l}|=0$ yielding a DSC of 0.
 Clinical Metrics:
@@
 For each scenario, we analyzed the performance of the loss functions across different groupings of MTV.
-3 Results
-3.1 Segmentation performance across different networks
-In this section, we present results for the segmentation performance of the three loss functions, Dice Loss (DL), Dice Focal Loss (DFL), and L1-weighted Dice Focal Loss (L1DFL) using the Attention U-Net and SegResNet architectures (Table 1).
-We report the average and median DSCs at the patient level, with mean values of true positive, false positive, and false negative counts on the test set and the F1 score performance.
-A one-tailed paired Wilcoxon signed-rank test was performed to evaluate the significance of the performance of L1DFL from those of DL and DFL separately, with corresponding p-values reported.
-Network Loss Function DSC Mean Rates F1 scores P values
-Mean Median TP FP FN
-Attention U-Net DL 0.54 $\pm$ 0.24 0.58 [0.44, 0.72] 0.81 $\pm$ 0.35 2.06 $\pm$ 1.89 0.19 $\pm$ 0.35 0.50 $\pm$ 0.28 $<0.01$*
-DFL 0.51 $\pm$ 0.26 0.54 [0.33, 0.73] 0.77 $\pm$ 0.36 2.65 $\pm$ 2.60 0.23 $\pm$ 0.36 0.44 $\pm$ 0.27 $<0.01$*
-L1DFL 0.58 $\pm$ 0.27 0.66 [0.51, 0.77] 0.77 $\pm$ 0.37 0.42 $\pm$ 0.65 0.23 $\pm$ 0.37 0.69 $\pm$ 0.34
-SegResNet DL 0.52 $\pm$ 0.30 0.60 [0.29, 0.76] 0.73 $\pm$ 0.39 0.73 $\pm$ 1.10 0.27 $\pm$ 0.39 0.62 $\pm$ 0.35 0.24
-DFL 0.52 $\pm$ 0.25 0.59 [0.41, 0.71] 0.81 $\pm$ 0.34 2.05 $\pm$ 2.00 0.19 $\pm$ 0.34 0.49 $\pm$ 0.27 0.015*
-L1DFL 0.57 $\pm$ 0.29 0.68 [0.47, 0.78] 0.76 $\pm$ 0.37 0.52 $\pm$ 0.76 0.24 $\pm$ 0.37 0.66 $\pm$ 0.34
-Table 1: Comparison of the loss functions on the test set based on mean and median patient-level DSC.
-The mean DSC, TP, FP, FN and F1 scores are presented with standard deviations and the median DSC with interquartile ranges.
-A detection is considered a true positive if the predicted mask overlaps with the voxel containing the SUVmax value in the ground truth.
-The P values reflect a one-tailed paired Wilcoxon signed-rank test comparing the median difference in DSC between L1DFL and DL and DFL, with * indicating statistical significance at the $\alpha=0.05$ level.
-The median DSCs are higher than the average DSCs, with SegResNet slightly outperforming Attention U-Net in DSC.
-However, Attention U-Net generally got better true positive rates except in the case of DFL, where SegResNet performed better with a $5\%$ increment.
-In both architectures, Dice Loss always performed better than Dice Focal Loss in minimizing false positives.
-However, both loss functions showed varied performances on the different networks regarding true positive rates.
-Specifically, while the Dice Loss surpassed the other loss functions with a true positive rate of 0.81 $\pm$ 0.35 on the Attention U-Net, Dice Focal Loss also achieved a similar performance but on the SegResNet architecture.
-Yet, this higher sensitivity of Dice Loss on Attention U-Net came at the cost of more false positives.
-Compared to a false positive rate of 0.73 $\pm$ 1.10 on SegResNet, the Dice Loss had a rate of 2.06 $\pm$ 1.89 on Attention U-Net, about three times the value on the former network.
-This imbalance in true and false positives on Attention U-Net led to a lower F1 score for Dice Loss despite its higher sensitivity.
-Conversely, L1DFL achieved the best balance of true positive detections and minimum false positives.
-For example, on Attention U-Net, where Dice Loss and Dice Focal Loss had mean false positive rates above 2.0, the L1DFL minimized the false detection rates.
-Although its FP rate slightly increased with SegResNet, its value was still low compared to DL and DFL.
-By lowering false positive rates while maintaining sensitivity to lesions, L1DFL yielded superior mean and median DSCs.
-Specifically, on Attention U-Net, L1DFL reached a median DSC of 0.66, outperforming DFL and DL by $22\%$ and $13\%$, respectively.
-Moreover, the high interquartile range of L1DFL indicates that its DSC values fell within a narrow range of high values.
-Although the one-tailed Wilcoxon signed-rank test comparing DL and L1DFL did not provide enough evidence to reject the null hypothesis of no significant difference between the two on the SegResNet architecture, statistically significant differences were obtained on the Attention U-Net.
-Also, comparing DFL versus L1DFL resulted in statistically significant differences in favor of L1DFL across both networks.
-Figure 3: Segmentation performance of the loss functions across two networks: Attention U-Net (top row) and SegResNet (bottom row).
-The first column shows the ground truth segmentation (GT) and the remaining columns represent the results for the loss functions: Dice Loss (DL), Dice Focal Loss (DFL), and L1-weighted Dice Focal Loss (L1DFL).
-The green-colored region denotes the ground truth annotation of the lesion, while the red-colored regions represent the segmentations produced by the models.
-The bounded area highlights the lesion in the ground truth image alongside the corresponding correct segmentation by the models.
-Red regions outside the bounded area indicate false positive segmentations.
-Figure 3 shows a sample segmentation performance of the loss functions on a test PET/CT image.
-The first-row segmentations are from Attention U-Net, while the second row is from SegResNet, with columns having different loss functions.
-The bounded region in the figure depicts correct segmentation according to the ground truth highlighting the differences in performance between the loss functions and the network architectures.
-In this example, Attention U-Net yielded better segmentation of the lesion with higher DSC than SegResNet.
-Yet, for both networks, the Dice and the Dice Focal Loss functions produced instances of false positives further impacting the overall DSC.
-L1DFL, however, remained robust, providing a consistent performance.
-3.2 Performance in single and multiple lesion scenarios
-We evaluated the performance of the loss functions in two different lesion scenarios: multiple lesions in one PET/CT volumetric image and only one lesion in the volume constituting the single lesion scenario.
-Figure 4 shows that the loss functions performed differently under these scenarios; a wider spread of DSC values is observed in the single lesion scenario than in the multiple scenario for both networks.
-Also, the DSCs beyond the 75th percentile for each loss function in the single lesion scenario were higher than those in the multiple lesion scenario, with more cases achieving DSCs greater than 0.8.
-Despite this observation, each loss function with Attention U-Net had higher median DSCs in the multiple-lesion scenario than in the single-lesion scenario (Figure 4a).
-Besides, in the single lesion case, all the loss functions have more lower outliers, reflecting occasional underperformance.
-On the other hand, for all the loss functions, there are fewer outliers in the multiple-lesion scenario.
-While similar observations are made on the SegResNet architecture, the median DSC of L1DFL in the single lesion scenario remained slightly higher than the median DSC in the multiple lesion case (Figure 4b).
-Figure 4: Box plots overlaid with swarm plots illustrating the distribution of Dice Similarity Coefficient (DSC) performance for the three loss functions across two lesion scenarios: M (lesion count $>2$) and S (lesion count $=1$).
-Each box plot’s top and bottom edges represent the interquartile range (IQR), while the horizontal lines within the boxes indicate the median DSC values.
-Whisker lengths are set to 1.5 times the IQR.
-(a) Performance on Attention U-Net architecture.
-(b) Performance on SegResNet.
-The number of cases in the single lesion scenario was N = 34 and for the multiple lesion scenario, the number of cases was N = 23.
-Dice Loss generally had higher median DSC for both architectures when compared with DFL, although there were variations in performance between lesion scenarios and architectures.
-In a single lesion scenario using the Attention U-Net architecture, for example, DL demonstrated more consistent performance as evidenced by an IQR of 0.33 compared to DFL with an IQR of 0.45.
-However, on SegResNet, DL’s performance became highly incoherent, with a wider IQR of 0.72 compared to the DFL’s narrower IQR of 0.37.
-In the case of multiple lesions, while both loss functions showed similar IQRs of 0.26 with SegResNet, on Attention U-Net, however, the DSC values from DL are more spread out with lower values in the lower quartile range compared to DFL and L1DFL.
-Thus, Dice Loss may not be consistent in some scenarios and architectures, particularly where the cases are more complex.
-In contrast, L1DFL outperformed DL and DFL in both lesion scenarios on the Attention U-Net.
-Its median DSCs on the multiple and single lesions were 0.67 and 0.65, respectively.
-These values are at least $11\%$ higher than the median DSCs of DL in both cases.
-Similarly, on SegResNet in the single lesion case, L1DFL achieved a median DSC of 0.69, while DL and DFL had 0.56 and 0.55, respectively.
-However, in the multiple lesion scenario, Dice Loss slightly outperformed L1DFL by $4\%$, although they both had similar IQR.
-3.3 Performance based on lesion molecular volume
-For each lesion scenario, we evaluated the performance of the loss functions across different TMTV thresholds at both the patient and lesion levels (Figure 5).
-Higher DSC values were generally observed when assessing the performance at the lesion level while the patient-level DSC values tended to be slightly lower due to false-positive segmentations.
-Notably, the Dice and Dice Focal Losses were more susceptible to false positives, resulting in a reduction of their median DSC by at least $10\%$ in the patient-level assessment.
-Figure 5: Patient and lesion-level median Dice Similarity Coefficient (DSC) performance of the different loss functions as a function of various ground truth tumor molecular volume thresholds.
-Each column shows the plots for the different lesion scenarios: single (S) and multiple (M).
-At both the patient and lesion levels, the line plots represent the trend of each loss function’s performance across different total molecular tumor volume (TMTV) and molecular tumor volume (MTV) thresholds, respectively, with thresholds $(t)$ based on the interquartile range from 0 to the 85th percentile.
-Median DSC is calculated for all lesions $(l)$ where volume $(v_{l}>t)$, and each color indicates a different loss function.
-The first row of plots (a) shows the performance of the loss functions with Attention U-Net.
-The second row (b) shows the performance of the loss functions with SegResNet.
-The third row (c) displays area plots of median tumor lesion activity (TLA) values according to the TMTV and MTV thresholds, where median TLA values are calculated for lesions with $(v_{l}>t)$.
-[…]
-[…]
-The Dice Loss performed better on SegResNet than on Attention U-Net, outperforming Dice Focal Loss at the patient level in both scenarios due to Dice Focal Loss’ higher false positive rate.
-However, both loss functions showed similar trends at the lesion level.
-The Dice Focal Loss on the other hand yielded more accurate segmentations with the Attention U-Net outperforming the Dice Loss at both scenarios.
-Nevertheless L1DFL maintained stable performance across both architectures and lesion scenarios.
-This consistency suggests the effectiveness of L1DFL’s dynamic weighting strategy, making it more robust to false positives at the patient level and achieving better overlap with the ground truth at the lesion level.
-3.4 Performance based on lesion dissemination
-To quantify the spatial distribution of lesions, we calculated $D_{\text{max}}$, the maximum Euclidean distance between any two lesions in the image.
-Figure 6: Plots illustrating the Dice Similarity Coefficients (DSC) across four interquartile ranges of $D_{\text{max}}$.
-The first row (a) shows the performance with Attention U-Net.
-The second row (b) shows the performance with SegResNet.
-The groups G0 (0-9 cm), G1 (9-11cm), G2 (11-14cm), and G3 (14-60cm) correspond to the interquartile ranges of $0-25\%$, $25-50\%$, $50-75\%$, and $75-100\%$, respectively.
-In the box plots, horizontal lines indicate the median, and whisker lengths are set to 1.5 times the interquartile range (IQR), with outliers represented as black dots.
-Accompanying bar plots (c) display the median total lesion activity (TLA) values for cases within each $D_{\text{max}}$ group.
-The different loss functions are color-coded for clarity.
-The total number of cases in each group was N = 6.
-For the performance assessment of the loss functions based on the $D_{\text{max}}$, we categorized the distances into the following groups: 0-9 cm (G0), 9-11 cm (G1), 11-14 cm (G2), and 14-60 cm (G3), representing the first, second, third and fourth quartile ranges, respectively.
-Patient-level analysis of DSC showed consistent improvement across these ranges (Figure 6).
-The figure also shows differences in performance across the different architectures.
-Generally, the Dice Loss performed better on SegResNet than on Attention U-Net but the Dice Focal Loss performed better on the latter architecture.
-The L1DFL on the other hand, yielded more consistent performance across the two networks, yet had a wider IQR on SegResNet than on Attention U-Net.
-On Attention U-Net, initially, median DSC values for all loss functions ranged from 0.4 to 0.6 (G0), increasing in the third quartile to 0.6-0.8 (G2).
-The results with SegResNet also show a similar trend; there is a general increase in DSC values up to the third quartile.
-For both networks, in the last quartile (G3), representing a more expansive lesion spread, all loss functions showed a drop in their performance metrics.
-Specifically, on Attention U-Net, the Dice Focal Loss performed the worst in this group, and L1DFL and Dice Loss presented similar performances, with a median of 0.62.
-Yet, the Dice Focal Loss had a smaller interquartile range, reflecting less spread across DSCs in that category.
-The median DSC of L1DFL dropped from 0.76 to 0.62 by $18.4\%$, and that of Dice Focal Loss dropped from 0.65 to 0.58 by $10.3\%$.
-Dice Loss also slightly declined in performance with a broader spread in IQR of its DSCs.
-On SegResNet, the Dice Loss had the worst performance in this group with L1DFL achieving the best segmentation accuracy.
-The performance of the loss functions in the lower $D_{\text{max}}$ ranges (G0 - G2) which characterize less lesion spread in the body further highlight network-specific variations.
-While the Dice Loss achieved similar high median DSC as L1DFL in these ranges with the SegResNet, it had the worst performance with Attention U-Net.
-Similarly, the SegResNet model trained with the Dice Focal Loss had a decline in obtaining accurate segmentations while pairing this loss function with Attention U-Net yielded competitive results.
-L1DFL remained robust to these architecture-specific variations consistently outperforming the baseline models.
-Figure 6 also depicts a higher median total lesion activity (TLA) across the different groups, indicating an increased lesion size or uptake value with greater lesion spread.
-Although model performance generally improves with higher lesion activity, the slight decrease in median DSC in the highest quartile of $D_{\text{max}}$, despite high TLA, suggests that other factors might contribute to the segmentation performance such as lesion site and count.
-4 Discussion
-In this work, we proposed and evaluated a novel loss function L1-weighted Dice Focal Loss, in segmenting metastatic prostate cancer lesions based on different lesion scenarios characterized by the number of lesions in an image.
-Specifically, we defined two scenarios: single lesion scenario which include whole body PET/CT images with only a single lesion present in the ground truth mask and multiple lesion scenario being images with more than one lesion.
-We compared the performance of our proposed loss function with the Dice and Dice Focal Losses.
-Moreover, we report how the size of a lesion and its spatial distribution influence the segmentation result.
-The dataset used in this study included tumors which had resurfaced after treatment with relatively low molecular volumes and standardized uptake values (SUV) thus introducing unique challenges for correct segmentations.
-Our findings showed that the performance of the loss functions is not uniform across different architectures.
-For instance, the results show that the Dice Loss generated more false positives on Attention U-Net compared to SegResNet, despite showing higher sensitivity on the former.
-Similarly, the Dice Focal Loss exhibited lower precision on the Attention U-Net architecture due to increased false positives.
-By evaluating performance at both the patient and lesion lesions, the effect of false positive rates was highlighted.
-There was a more than $10\%$ decrease in the median DSC values at the lesion level for both loss functions when assessments were made at the patient level.
-L1DFL, in contrast, was more robust to these false detections.
-[…]
-Furthermore, comparing the results obtained with SegResNet on the single lesion scenario to the multiple lesion scenario points to the influence of additional factors to segmentation accuracy other than tumor volume and activity.
-Due to the summation of the TLA values of individual lesions in the multiple lesion scenario, higher activity values are obtained than in the single lesion case yet, lower DSC values are observed.
-This observation could suggest the introduction of further complexities as lesion count increases.
-In fact, with more lesions, different anatomical sites may be involved, and the effect of other organs with high uptake values like the bladder could be pronounced [31].
-Besides, the evaluation based on lesion spread, $D_{\text{max}}$, highlights the decrease in the models’ performance as distribution across the body becomes more expansive.
-The Dice Loss is based on overlap measurement and it inherently balances sensitivity with specificity, making it more robust in scenarios with significant data imbalance such as in lesion segmentation tasks [16].
-Therefore, distribution-based losses are often combined with the Dice Loss or its variants to improve their robustness.
-However, without adapting its focus to classification difficulty, the Dice Loss fails to adjust dynamically toward heterogeneous lesion characteristics.
-Despite the unique challenges posed by the tumor heterogeneity, the L1-weighted Dice Focal Loss generally yielded better segmentation performance, achieving better balance in false detection rates.
-This loss function comprises a dynamic weighting strategy that adapts the model’s focus based on sample difficulty.
-For easier-to-classify examples (characterized by predicted probabilities $P>0.5$), L1DFL applies a reduced penalty while penalizing more challenging examples ($P<0.5$) more heavily (Figure 7).
-Figure 7: Plot of loss versus probability for samples with the ground truth label as 1.
-The curves represent the behavior of each loss function, highlighting how they respond to varying probabilities of correctly classifying examples.
-The plot illustrates the differences in how each loss function handles both easy-to-classify and hard-to-classify cases based on their predicted probabilities.
-Similarly, the Dice Focal Loss increases the penalty on samples with a lower probability of correctly classifying the ground truth label, P(G), encouraging the model to focus more on ”hard” samples, but it applies a much steeper penalty.
-Thus, it tends to lose the Dice Loss component, where positive and negative examples remain equally weighted, potentially leading to more false positive rates due to misclassification of background areas as lesions [18].
-On the other hand, the weighting strategy of L1DFL allows for a more balanced approach to penalizing misclassification, consequently reducing false positive rates.
-Our results indicate that L1DFL yielded the highest F1 and DSCs in this study, underlining its more robust performance across various lesion scenarios and levels of lesion dissemination.
-This adaptive weighting strategy shows promise in improving segmentation performance in complex metastatic imaging cases.
-5 Conclusion
-We evaluated the performance of a custom loss function, L1-weighted Dice Focal Loss (L1DFL), in the detection and segmentation of metastatic prostate cancer lesions in PSMA PET/CT scans across single and multiple lesion scenarios, characterized by lesion count.
-L1DFL leverages a dynamic weighting scheme that is based on the L1 norms of the predicted probabilities and ground truth labels to modulate the Dice Loss and then combined with the Focal Loss to emphasize challenging pixel classifications.
-We compared L1DFL’s performance against Dice Loss and Dice Focal Loss using two 3D CNN architectures, Attention U-Net, and SegResNet.
-While Dice Loss demonstrated sensitivity for small lesions, it usually suffered from low precision.
-Dice Focal Loss showed higher false positive rates and variable performance across the different networks.
-On the other hand, L1DFL consistently achieved a balanced segmentation performance across the lesion scenarios and architectures, achieving superior F1 scores and Dice Similarity Coefficients.
-Future work will be directed at exploring the generalizability of L1DFL to other medical imaging datasets and tasks.
-6 Acknowledgments
-This work was supported by the Canadian Institutes of Health Research (CIHR) Project Grants PJT-162216 and PJT-173231.
-We also acknowledge computational resources and services provided by Microsoft AI for Health.